### v2

# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

Abstract
In this article, we provide a stochastic sampling algorithm with polynomial complexity in fixed dimension that leverages the recent advances on diffusion models where it is shown that under mild conditions, sampling can be achieved via an accurate estimation of intermediate scores across the marginals $(p_{t})_{t\geq 0}$ of the standard Ornstein-Uhlenbeck process started at $\mu$, the density we wish to sample from. The heart of our method consists into approaching these scores via a computationally cheap estimator and relating the variance of this estimator to the smoothness properties of the forward process. Under the assumption that the density to sample from is $L$-log-smooth and that the forward process is semi-log-concave: $-\nabla^{2}\log(p_{t})\succeq-\beta I_{d}$ for some $\beta\geq 0$, we prove that our algorithm achieves an expected $\epsilon$ error in KL divergence in $O(d^{7}(L+\beta)^{2}L^{d+2}\epsilon^{-2(d+3)}(d+m_{2}(\mu))^{2(d+1)})$ time with $m_{2}(\mu)$ the second order moment of $\mu$. In particular, our result allows to fully transfer the problem of sampling from a log-smooth distribution into a regularity estimate problem. As an application, we derive an exponential complexity improvement for the problem of sampling from an $L$-log-smooth distribution that is $\alpha$-strongly log-concave outside some ball of radius $R$: after proving that such distributions verify the semi-log-concavity assumption, a result which might be of independent interest, we recover a $poly(R,L,\alpha^{-1},\epsilon^{-1})$ complexity in fixed dimension which exponentially improves upon the previously known $poly(e^{LR^{2}},L,\alpha^{-1},\log(\epsilon^{-1}))$ complexity in the low precision regime.
1 Introduction
Given some potential $V:\mathbb{R}^{d}\to\mathbb{R}$ such that $\int e^{-V(x)}\mathrm{d}x<\infty$, the sampling problem consists into obtaining a sample from some distribution $p$ such that $p$ is $\epsilon$-close to $\mu\propto e^{-V}$ with respect to some divergence while maintaining the complexity, a.k.a. the number of queries to $V$ and possibly to its derivatives, as low as possible with respect to $\epsilon$ and the other parameters of the problem. As in Euclidean optimization, appealing complexity bounds can be obtained when $\mu$ is assumed to be $L$-log-smooth and $\alpha$-strongly log-concave or equivalently, when the potential $V$ is assumed to have an $L$-Lispchitz gradient and to be $\alpha$-strongly convex. Namely, denoting the condition number $\kappa=L/\alpha$, it has been recently shown that the Underdamped Langevin Algorithm (ULA) [14, 49], one of the most popular sampling algorithm, could provide a sample $X\sim p$ such that $\mathop{\mathrm{KL}}\nolimits(p,\mu)\lesssim\epsilon$ in $\tilde{O}(d^{1/3}\kappa\epsilon^{-2/3})$ time under minor modifications [43, 20, 7]. Similarly, it was shown that the Metropolis Adjusted Langevin Algorithm (MALA) [48, 3], a debiased version of Langevin, could achieve a complexity of $\tilde{O}(d^{1/2}\kappa\log^{4}(1/\epsilon))$ for a well-chosen warm start [2]; in both cases, the dependence is sub-linear in the dimension. However, because strong log-concavity implies unimodality, this framework is far from satisfactory when it comes to obtaining guarantees for real-world applications, where the distributions to sample from can be highly multimodal [30, 17]. One natural relaxation is to confine mutlimodality within some fixed ball, also known as the strong dissipativity assumption [13, 31, 32, 19] where we assume that the distribution is strongly log-concave outside this ball. Note that this assumption is much more realistic as it encompasses several types of Gaussian mixtures and other popular Bayesian mixture models [34, 33, 19]. Yet, this extra flexibility currently comes at a heavy price as known complexity bounds now degrade exponentially with the size of the non-convex region: if we assume that $\mu$ is $\alpha$-strongly log-concave outside the centered ball of radius $R$, ULA achieves a complexity of $\tilde{O}(e^{32LR^{2}}\kappa^{2}\frac{d}{\epsilon^{2}})$ and MALA achieves a complexity of $O(\frac{e^{40LR^{2}}}{\alpha}\kappa^{3/2}d^{1/2}(d\log(\kappa)+\log(1/\epsilon))^{3/2})$ [31]. Note that even if those are indeed worst case bounds, it has been well observed in practice that the aforementioned samplers suffer from metastability: as the distance between the modes grows, samples get stuck in one local mode, leaving the other ones unexplored [44, 26] (see also Chehab et al., [9] for formal lower bounds in continuous time). While there is no hope for a general polynomial bound as it was shown in Lee et al., [29] that even sampling from a $2$-modes distribution could require up to $2^{cd}$ queries to $V$ with $c$ a universal constant, sampling from a mutlimodal density in low dimension with provable polynomial bounds is still an open problem. This raises the following informal question:
“In fixed dimension, can we sample from a multimodal distribution in polynomial time?”
1.1 Related work
Sampling from multimodal distributions has been an active area of research for several decades now [37, 15, 38]. Indeed, it has long been observed that Monte Carlo Markov Chain (MCMC) based on local exploration of the space, such as ULA and MALA, are prawn to metastability and remain stuck in local modes [47, 21, 42]. We review in this paragraph the main alternatives to Langevin based samplers that have been developed and the guarantees they provide.
Histogram methods
Given some compact set $S\subset\mathbb{R}^{d}$, histogram methods build a regular $n$-grid over $S$ and sample a point $s_{i}$ of the grid according to the probability $e^{-s_{i}}/\sum_{i=1}^{n}e^{-V(s_{j})}$. If $V$ is assumed to be $\alpha$-strongly convex and $L$-smooth for instance, then, if $S$ is taken to be a ball of radius $r$ centered at $x^{*}$ the global minimum of $V$, the typical error of this method scales in $O(L(r/n)^{1/d}r+e^{-\alpha r^{2}})\,,$ thus inducing a typical complexity of $\tilde{O}((L/\epsilon)^{d})$ which is indeed polynomial in fixed dimension. However, if instead of being centered at the global minimum $x^{*}$, the ball is centered at some point $x$, the error remains $O(1)$ as long as $r\leq\|x-x^{*}\|$ as only the tail of $\mu$ is now discretized. When $V$ is smooth and strongly convex, $x^{*}$ can be approximated up to arbitrary precision yet in the case where $V$ is assumed to be strongly convex only outside of some unknown ball, one cannot localize the global minimum $x^{*}$ even in one dimension. In other terms, unless we have prior knowledge on the localization of the mass of $\mu$, histogram methods can induce an arbitrarily large error.
Importance Sampling (IS)
Instead of focusing on the problem of sampling from $\mu$, IS focuses on the related, yet slightly simpler problem of approximation of integrals under $\mu$: given some bounded function $f$, IS methods approximate the integral $\mu(f)\coloneqq\int_{x}f(x)\mathrm{d}\mu(x)\,.$ For any $W$ such that $\int e^{-W(x)}\mathrm{d}x<+\infty$, denoting $\nu\propto e^{-W}$, the integral above can be re-written as an expectation under $\nu$ as $\mu(f)=\mathbb{E}_{Y\sim\nu}[f(Y)\frac{\mathrm{d}\mu}{\mathrm{d}\nu}(Y)]\,.$ In particular, if we can access $n$ i.i.d. samples $(x_{i})_{1\leq i\leq n}$ from the proposal $\nu$, this quantity can be approximated as $\hat{\mu}(f)=\frac{\sum_{i=1}^{n}f(x_{i})e^{-V(x_{i})+W(x_{i})}}{\sum_{i=1}^{n}e^{-V(x_{i})+W(x_{i})}}\,.$ It was shown in Robins et al., [41] that the average quadratic error of this estimator scales as $O\biggl{(}\frac{\chi^{2}(\mu,\nu)}{n}\biggr{)}\,.$ Yet, unless $\nu$ is carefully chosen, this bound may blow up to infinity even for simple densities: if $\mu\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $\nu\sim\mathcal{N}(\mu_{2},\sigma_{2}^{2})$, it holds that $\chi^{2}(\mu,\nu)=\frac{1}{2}\biggl{(}\frac{\sigma_{2}^{2}}{\sigma_{1}\sqrt{2\sigma_{2}^{2}-\sigma_{1}^{2}}}e^{\frac{(\mu_{1}-\mu_{2})^{2}}{2\sigma_{2}^{2}-\sigma_{1}^{2}}}-1\biggr{)}\,.$ In particular, not only this bound blows up when the variance of $\nu$ is not at least half the variance of $\mu$ but also, this bound always degrades exponentially with the quadratic distance between the mode of the proposal and the mode of the target $\mu$. Hence, even in this one dimensional strongly log-concave toy example, IS requires a precise a priori knowledge of $\mu$ to tailor a good proposal $\nu$ (and also requires to be able to sample from $\nu$). Nevertheless, we discuss in the last paragraph of this section the recent work of Chak, [8] who shows that if the localization of the mass of $\mu$ is known, as in histogram methods, a suitable proposal $\nu$ can be designed with appealing bounds.
Tempering heuristics
Instead of directly sampling from $\mu\propto e^{-V}$, we start to sample from a flattened version of $\mu$ given by $\mu^{\beta}\propto e^{-\beta V}$ for $\beta$ small and we progressively increase $\beta$ to $1$. When $\mu$ is assumed to be a finite mixture of the same shifted $\alpha$-strongly log-concave and $L$-log-smooth distribution, Lee et al., [29] proved that for a well-chosen (stochastic) sequence of flattened distributions $\mu^{\beta_{i}}$, sampling up to precision $\epsilon$ in $\mathop{\mathrm{KL}}\nolimits$ can be achieved in $poly(L,\alpha^{-1},w_{min},d,\epsilon^{-1},R)$ time with $R$ the diameter of the shifts and $w_{\min}$ the minimum weight in the mixture. Yet, despite the strong guarantees it offers, this framework excludes many practical cases such as mixtures of Gaussians with different variances for instance: in that case, the tempering strategy even fails in low dimension as it attributes biased weights to the different modes.
Reverse Diffusion (RD) methods
As we shall review in detail in Section 2, recent works show that, under milder and milder assumptions [12, 10, 16, 5], the problem of sampling can be transferred into the problem of approximating the score in $L^{2}$ distance along the standard Ornstein-Uhlenbeck (OU) process initialized at the targeted distribution $\mu$. Over the past year, several works sought to leverage this result and focused on the task of estimating these intermediate scores given an oracle access to $V$ and eventually to its gradient. Their guarantees can be summarized as follows: in Ding et al., [18], denoting $\pi^{\eta,\Sigma}$ the $d$-dimensional Gaussian density with mean $\eta$ and covariance matrix $\Sigma$, the authors proved that if there existed known $\eta,\Sigma$ such that $\mathrm{d}\mu/\mathrm{d}\pi^{\eta,\Sigma}$ is globally $L$-Lipschitz and strictly bounded from below by $\xi$, then their algorithm achieved a complexity $poly(L,\xi^{-1},d,\epsilon^{-1})$. We show in Appendix C that even for the simple mixture $\mu=\frac{1}{2}\mathcal{N}(1,1)+\frac{1}{2}\mathcal{N}(-1,1)\,,$ the previous conditions do not hold. In Huang et al., [25], it is shown that under the assumption that the OU process remains $L$-log-smooth along the trajectory, denoting $m_{2}(\mu)=\int\|x\|^{2}\mathrm{d}\mu(x)$ the second moment of $\mu$, their algorithm could reach a complexity of $e^{O(L^{3}\log^{3}((Ld+m_{2}(\mu))/\epsilon))}$. In He et al., [22], it is shown that under a sub-quadratic growth assumption on $V$ (which is weaker than log-smoothness), i.e. there exists $L>0$ such that $V(x)-V(x^{*})\leq L\|x-x^{*}\|^{2}\,,$ with $x^{*}$ a global minimizer of $V$, their algorithm could reach at best a complexity of $O(L^{d/2}\epsilon^{-d}e^{(L\|x^{*}\|^{2}+\|x_{N}\|^{2})})\,,$ where $x_{N}$ is the sample output by their algorithm. Under the reasonable (and desirable) assumption that $\mathbb{E}[||x_{N}||^{2}]\asymp m_{2}(\mu)$, Jensen inequality yields that their complexity scales at best as $O(L^{d/2}\epsilon^{-d}e^{(L\|x^{*}\|^{2}+m_{2}(\mu))})\,.$ Hence, we observe that previous approaches based on reverse diffusion suffer at best from exponential complexity either with respect to $L$ or to $m_{2}(\mu)$.
1.2 Our contributions
Polynomial time sampling under semi-log-concavity of the Ornstein-Uhlenbeck (OU) process
As in the aforementioned work, we rely on the recent advances on reverse diffusion and focus on the task of estimating the intermediate scores $\nabla\log(p_{t})$ with $(p_{t})_{t\geq 0}$ the marginals of the standard OU process initialized at $\mu$. Yet, unlike all previous works, we provide fully polynomial time guarantees in fixed dimension when $\mu$ is assumed to be log-smooth and when the forward process has semi-log-concave marginals.
Theorem 1 (Main result, informal)
Denoting $p_{t}^{V}$ the density at time $t$ of the forward OU process started at $\mu\propto e^{-V}$, then under the assumption that $V$ is $C^{2}$ and $L$-smooth, that $\mu$ has finite second moment $m_{2}(\mu)$ and that $-\nabla^{2}\log(p_{t}^{V}),-\nabla^{2}\log(p_{t}^{2V})\succeq-\beta I_{d}$ uniformly over $\mathbb{R}^{d}$, then there exists a stochastic algorithm that outputs a sample $X\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon$ in $O(d^{7}L^{d+2}\epsilon^{-2(d+3)}(L+\beta)^{2}(m_{2}(\mu)+d)^{2(d+1)})$ time where $\lesssim$ hides a universal constant as well as log quantities in $d,\epsilon^{-1},L,\beta,m_{2}(\mu)$.
Hence, when the dimension $d$ is fixed, we obtain a polynomial time guarantee. Note that this result pushes one step further the works of Chen et al., 2023b [12], Chen et al., 2023a [10], Conforti et al., [16], Benton et al., [5] who showed that the sampling problem could be transferred into a score approximation problem yet leaving this crucial question open. Using our estimator of the intermediate scores, we show that in fixed dimension, sampling can be framed as a regularity estimate problem: for any class of smooth potentials $V$ such that $\nabla^{2}\log(p_{t}^{V}),\nabla^{2}\log(p_{t}^{2V})$ are bounded from above, we can sample from $\mu\propto e^{-V}$ in polynomial time.
Polynomial time sampling for smooth and strongly dissipative distributions
Deriving classes of potentials $V$ that allow quantitative bounds on $\nabla^{2}\log(p_{t}^{V})$ is an area of research of its own as it allows to recover transport maps from the standard Gaussian to measures of the form $\mu\propto e^{-V}$ with quantitative Lipschitz bounds. Such bounds were originally proven for the class of smooth and strongly convex potentials [27] and later derived for the classes of compactly supported distributions and mixtures of Gaussians with the same variance respectively [35]. In this work, we show that such bounds still hold for smooth potentials that are only assumed to be strongly convex outside some ball. Using Theorem 1, this allows us to recover polynomial complexity bounds for sampling from this class of distributions.
Theorem 2 (Application to smooth and strongly dissipative potentials)
Assume that $V$ is $C^{2}$, $L$-smooth and $\alpha$-strongly convex outside some ball of radius $R$, then there exists a stochastic algorithm that that returns a sample $X\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon$ in $O(d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)})$ time.
As previously mentioned, this bound exponentially improves upon the $O(e^{32RL^{2}}d\kappa^{2}\epsilon^{-2})$ complexity of ULA [31] and the $O(\frac{e^{40RL^{2}}}{\alpha}\kappa^{3/2}d^{1/2}(d\log(\kappa)+\log(1/\epsilon))^{3/2})$ complexity of MALA [31] in the low dimension and low precision regime. We also show in Section 4 that our results exponentially improve upon Huang et al., [25] and He et al., [22].
1.3 Discussion on concurrent work
In the very recent paper Chak, [8], polynomial time guarantees were also derived for the task of sampling from a distribution that is log-smooth and strongly log-concave outside of a ball that is centered at 00, yet with better exponents than ours. To achieve this, the author relies on IS with a carefully chosen proposal distribution $\nu$: a cutoff parameter $M$ is chosen such that the proposal distribution $\nu$ matches with $\mu\propto e^{-V}$ whenever $V(x)\geq M$ and is flat everywhere else. In Theorem 2.3, such a proposal $\nu$ with matching tails is proven to be a suitable choice for IS. Furthermore, when $M$ is sufficiently large, it is shown in Proposition 5.1 therein that the proposal $\nu$ is log-concave and can thus be safely sampled from using ULA or MALA for instance. However, the cutoff value of $M$ that ensures log-concavity requires estimates on the log-smoothness $L$ and the radius $R$ which may not be available in practice: we discuss in Appendix C a simple, one dimensional example where $R$ depends on the localization of the modes which are unknown a priori. In contrast, our method does not require the knowledge of any of these constants. Furthermore, our method still provides quantitative guarantees under Assumption 1 which is more general than strong log-concavity outside a ball as shown in Section 4.
2 Preliminaries
In this section, we introduce the reverse diffusion framework and review the main existing bounds on how the resulting sampling error is affected by the error on the intermediate scores. Then, we show that these intermediate scores can be re-written as a ratio of expectations and detail how this re-writing differs from previous works.
2.1 Reverse diffusion: from sampling to score estimation
One of the core tool for sampling from $\mu\propto e^{-V}$ is the Langevin diffusion process $\mathrm{d}X_{t}=–\nabla V(X_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}\,,$ where $B_{t}$ is the standard Brownian motion. When discretized with a Euler-Maruyama scheme, this forward equation becomes the well-known Underdamped Langevin Algorithm (ULA, [40]) which is readily implementable whenever we can access $\nabla V$. One of the fundamental questions to determine the quality of the sample output by ULA is the speed of convergence of Langevin towards the equilibrium $\mu$. When $\mu$ is $\alpha$-strongly log-concave, and more generally when $\mu$ verifies an $\alpha$-log-Sobolev inequality [4], exponential convergence in KL occurs at an $\alpha$ rate [46]. Yet, if $\mu$ is multimodal, strong log-concavity no longer holds and the log-sobolev constant $\alpha$ is either no longer defined either scales exponentially with the distance between the modes [11, 9]. This exponential dependence is in correspondence with a phenomenon known as metastability where the process $X_{t}$ gets stuck in a local mode of $\mu$. Reverse diffusion methods emerged as an alternative to Langevin based samplers in order to overcome metastability and were first introduced to the ML community in Song and Ermon, [45]. Instead of targeting $\mu$, one targets the standard Gaussian measure with a Langevin diffusion initialized at $\mu$. We thus obtain the so-called forward process
$\begin{cases}&\mathrm{d}X_{t}=-X_{t}\mathrm{d}t+\mathrm{d}B_{t}\,,\\ &X_{0}\sim\mu\,,\end{cases}$ (1)
which corresponds to the standard Ornstein-Uhlenbeck (OU) process initialized at $\mu$. Note that since we are now targeting the standard Gaussian which is $1$-strongly log-concave, the resulting process converges exponentially fast to the equilibrium. In order to sample from $\mu$ we consider the semi-discretized backward process: given a horizon $T$ that we discretize as $0=t_{0}\leq t_{1}\leq\cdots t_{N-1}\leq t_{N}=T$, the discretized process reads
$\mathrm{d}Y_{t}=Y_{t}\mathrm{d}t+2\nabla\log(p_{t_{N-k}})(Y_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}\,,~{}~{}t\in]T-t_{N-k},T-t_{N-(k+1)}]\,,$ (2)
with $Y_{0}\sim p_{T}$ and where $p_{t}$ is the distribution of the forward process Eq. 1 at time $t$. Note that this reverse process cannot be readily implemented for two reasons: first, it requires the knowledge of the intermediate scores $\nabla\log(p_{t_{k}})$ which are not available in closed form. Second, it requires sampling from the distribution $p_{T}$. Nevertheless, if one can access a proxy $s_{t_{k}}$ of the scores $\nabla\log(p_{t_{k}})$, and considering $T$ large enough so that $p_{T}\approx\pi$, we can implement instead
$\mathrm{d}Y_{t}=Y_{t}\mathrm{d}t+2s_{t_{k}}(Y_{k})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}\,,~{}~{}t\in]T-t_{N-k},T-t_{N-(k+1)}]\,,$ (3)
with $Y_{0}\sim\pi$ and where all iterations can be solved in closed form. Because the forward process Eq. 1 converges exponentially fast, we can expect the initialization error $Y_{0}\sim\pi$ instead of $Y_{0}\sim p_{T}$ to be small for a large time $T$. Furthermore, if the proxies $s_{t_{k}}$ are sufficiently accurate, one can expect that the process output by the approximate scheme Eq. 3 has a distribution that is close to the target $\mu$. Over the past three years, several works provided quantitative bounds of the error induced by the discretization, the use of an approximate score and the initialization error with respect to different divergences and under various assumptions [6, 28, 12, 10, 16]. Yet, we shall rely exclusively on the following theorem as they are the most suited to our framework.
Theorem 3 (Chen et al., 2023a [10])
Assume that $\mu$ is $L$-log-smooth and has finite second order moment $m_{2}(\mu)=\int\|x\|^{2}\mathrm{d}\mu(x)$ and choose the following (backward) discretization: $t_{k-1}=t_{k}-c\min(\max(t_{k},1/L),1)$ for $k\geq 2$ with $c=(\log(L)+T)/N$. Then, if $c$ verifies $c\leq\frac{1}{2d}$, denoting $p$ the distribution of the sample $Y_{T}$ output by Eq. 3, it holds that
$\mathop{\mathrm{KL}}\nolimits(\mu,p)\lesssim(d+m_{2}(\mu))e^{-T}+\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}+\frac{d^{2}(\log(L)+T)^{2}}{N}\,,$
where $\lesssim$ hides a universal constant. If we make the stronger assumption that $p_{t}^{V}$ remains $L$-log smooth along the trajectory, then it holds using the constant step size $t_{k}=kT/N$ that
$\mathop{\mathrm{KL}}\nolimits(\mu,p)\lesssim(d+m_{2}(\mu))e^{-T}+\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}+\frac{dT^{2}L^{2}}{N}\,,$
where $\lesssim$ hides a universal constant.
In the case where we have an a priori bound on the log-smoothness of $\mu$, we can use the first statement of the theorem that yields a tighter bound in fixed dimension throught the use of a varying step-size depending on $L$. If no such prior knowledge is available, we can use a constant step size at the cost of a stronger assumption and a weaker bound.
Remark 4
In its original statement, Chen et al., 2023a [10, Theorem 5] assumes that
$\frac{1}{T}\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}\leq\epsilon^{2}\,,$
and directly provides the upper bound with the $T\epsilon^{2}$ term instead. Yet, an inspection of their Proposition 8 shows that we can safely use $\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}$. Similarly, it assumes that $c$ should be lower than $\frac{1}{Kd}$ with $K$ some universal constant. Yet a careful inspection of Lemma 13 and Lemma 17 shows that we can take $K=2$.
The previous theorems show that under mild assumptions that notably allow for multimodality, the problem of sampling from $\mu$ can be transferred into a score approximation problem along the forward process. In the next subsection, we build an estimator for these intermediate scores that is tractable given the knowledge of the unnormalized density $\mu\propto e^{-V}$, and show how it differs from previous works.
2.2 Construction of our estimator
The key observation to derive an estimator of the intermediate scores is that the forward process Eq. 1 is nearly available in closed form. Indeed, Eq. 1 integrates to $X_{t}=e^{-t}X_{0}+B_{1-e^{-2t}}\,,$ where $B_{s}$ is the standard $d$-dimensional Brownian motion evaluated at time $s$. In particular, conditionally on $X_{0}$, $X_{t}$ has a normal distribution $X_{t}|X_{0}\sim\mathcal{N}(e^{-t}X_{0},(1-e^{-2t})I_{d})$. By Bayes rule, the distribution $p_{t}^{V}$ of the forward process $X_{t}$ initialized at $\mu\propto e^{-V}$ thus integrates up normalizing constants to $p_{t}^{V}(z)\propto\int e^{-\frac{\|e^{-t}x-z\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}\mathrm{d}x\,.$ We thus recover that the score of the forward process re-writes as $\nabla\log(p_{t}^{V})(z)=\ \mathbb{E}_{Y\sim q_{t,z}}\biggl{[}\frac{z-e^{-t}Y}{1-e^{-2t}}\biggr{]}\,,$ where we denoted $q_{t,z}(x)\propto e^{-\frac{\|e^{-t}x-z\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}$. As noted in Huang et al., [24], Huang et al., [25] and He et al., [22], if samples $y_{i}\sim q_{t,z}$ are available, one can recover an efficient empirical estimator of the score given by $\frac{-1}{n(1-e^{-2t})}\sum_{i=1}^{n}y_{i}$ ready to be plugged in the reverse diffusion process Eq. 3. However, by doing so, the problem of sampling from $e^{-V}$ is simply transferred to the problem of sampling from $q_{t,z}$, which becomes almost as hard as $t$ grows. In Huang et al., [24], the intermediate sampling problems are solved using the ULA algorithm and in He et al., [22], the authors use a rejection sampling scheme which is similar in spirit to importance sampling. Hence, it is not surprising that their final upper-bounds do not improve over standard methods as they employ them to solve intermediate problems that are progressively as hard as the original problem. Instead, we observe that after applying the change of variable $y_{t}=ze^{-t}-x$ both on the numerator and on the denominator, we obtain a score formula which re-writes as a ratio of expectations under standard Gaussians $\nabla\log(p_{t}^{V})(z)=\frac{-1}{1-e^{-2t}}\frac{\mathbb{E}[Y_{t}e^{-V(e^{t}(z-Y_{t}))}]}{\mathbb{E}[e^{-V(e^{t}(z-Y_{t}))}]}\,,$ with $Y_{t}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$. Unlike the previous identity, which involves an expectation under $q_{t,z}$, the expectations above are computationally cheap as we can easily sample from $Y_{t}$. While conventional statistical wisdom would suggest both the numerator and the denominator to be estimated with independent samples, we voluntarily choose to correlate them and implement instead
$\hat{s}_{t,n}(z)\coloneqq\frac{-1}{1-e^{-2t}}\frac{\sum_{i=1}^{n}y_{i}e^{-V(e^{t}(z-y_{i}))}}{\sum_{i=1}^{n}e^{-V(e^{t}(z-y_{i}))}}\,,$ (4)
where the $y_{i}$ are independent Gaussians such that $y_{i}\sim\mathcal{N}(0,(1-e^{1-2t})I_{d})$ ; we shall refer to this estimator as self-normalized as common in the sampling literature [1]. Depending on the locations of the modes of $V$ we expect that this estimator behaves reasonably well in some regions of the space and poorly in others. Yet, the key property of self-normalized estimators is that they remain nearly bounded: in our case, it holds uniformly in $z,t$ that $\|\hat{s}_{t,n}(z)\|\leq\frac{\max_{i}\|y_{i}\|}{1-e^{-2t}}\sim\sqrt{d\log(n)}\,.$ This boundedness will allow us to derive a non-asymptotical control on the quadratic error that we present in the next section.
Remark 5
The estimator used in Ding et al., [18] is very similar to ours and simply differs by the fact that the authors discretize a Föllmer flow which corresponds to a reparametrized in time version of the standard OU process and where an arbitrary Gaussian may be targeted. The authors obtain strong guarantees yet under assumptions that exclude almost all distributions (see Appendix C). As a consequence, the authors avoid most of the technical difficulties we were confronted with.
3 Proof of the main result
In this section, we prove a non-asymptotic bound on the variance of the estimator presented above. First, we derive a general non-asymptotic bound that does not rely on any particular assumptions. Using the estimates provided in Mikulincer and Shenfeld, [36] and our additional semi-log-concavity assumption, this upper bound becomes fully explicit in the zeroth and second order moments of the ratio $\Phi_{t}=\frac{p_{t}^{2V}}{p_{t}^{V}}$. We then provide upper-bounds for these moments that again heavily leverage the estimates of Mikulincer and Shenfeld, [36]. Finally, we combine our results with those of Chen et al., 2023a [10] to state our main theorem.
3.1 Non asymptotical upper-bound of the variance
Before studying our ratio estimator defined in Eq. 4, we start by separately upper-bounding the variances of its numerator and denominator as a function of the forward process.
Proposition 6 (Variance of estimators)
Let $y_{1},\dots,y_{n}$ i.i.d. distributed as $\mathcal{N}(0,(1-e^{-2t})I_{d})$ and denote $\hat{N}(z)$ and $\hat{D}(z)$ the numerator and denominator of the estimator defined in Eq. 4, we have:
$\mathbb{E}[\|\hat{N}(z)-N(z)\|^{2}]$ $\leq\frac{p_{t}^{2V}Z_{2V}e^{-td}}{n}\left(\Delta\log(p_{t}^{2V})(z)-\frac{\Delta\log(\pi)(z)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}(z)\right)\,,$
$\mathbb{E}[\|\hat{D}(z)-D(z)\|^{2}]$ $\leq\frac{p_{t}^{2V}(z)Z_{2V}e^{-td}}{n}\,.$
The proof is deferred to Appendix A.2. We can now derive a non asymptotic quadratic error for our estimator.
Proposition 7 (Non-asymptotic bound of the quadratic error)
For all $z\in\mathbb{R}^{d}$, $n$ and $t>0$, it holds that
$\mathbb{E}\left[\|\hat{s}_{t,n}(z)-\nabla\log(p_{t}^{V})(z)\|^{2}\right]\leq\frac{32e^{2}(d+\log(n))}{n(1-e^{-2t})}\left(1+\theta_{t}(z)\frac{p_{t}^{2V}(z)Z_{2V}e^{td}}{(p_{t}^{V}(z)Z_{V})^{2}}\right)\,,$
with $\theta_{t}(z)=\Delta\log(p_{t}^{2V})(z)-\frac{\Delta\log(\pi)(z)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})(z)\|^{2}+\|\nabla\log(p_{t}^{2V})(z)\|^{2}+1$.
The complete proof is left to Appendix A.3 yet we briefly sketch the main arguments. We split the expectation on the event $A$ where the empirical denominator $\hat{D}$ (resp. the empirical numerator $\hat{N}$) is not too small (resp. not too large) with respect to its expectation $D$ (resp. $\|N\|$) and on the complementary $\bar{A}$. Across $A$, we use a Taylor expansion of order $2$ to make the variance of the numerator and of the denominator appear and we then use the previous proposition. Conversely, the estimator remains almost bounded on $\bar{A}$. We use Tchebychev inequality to upper-bound $\mathbb{P}(\bar{A})$ and make the variances of the numerator and of the denominator appear again, which concludes the proof.
Remark 8
Generic results on self-normalized estimators were derived in Agapiou et al., [1, Theorem 2.3]. Yet, it would have involved an additional $1/(p_{t}^{V})^{4}$ term which would have caused the integrated error $\int\mathbb{E}\left[\|\hat{s}_{t,n}(z)-\nabla\log(p_{t}^{V})(z)\|^{2}\right]\mathrm{d}p_{t}^{V}(z)$ to diverge.
Hence, we observe that the average quadratic error is related to the properties of the forward process Eq. 1 started at $\mu\propto e^{-V}$ and $\mu^{2}\propto e^{-2V}$ respectively. In particular, it involves the Laplacian of $\log(p_{t}^{2V})$ as well as the gradients of $\log(p_{t}^{V})$ and $\log(p_{t}^{2V})$. Hence, a first natural assumption is to bound the hessians of $\log(p_{t}^{V})$ and $\log(p_{t}^{2V})$ from above along the trajectory, a.k.a. semi-log-concavity along the forward.
Assumption 1 (Semi-log-concavity of the forward processes)
There exists $\beta\geq 0$ such that
$-\nabla^{2}\log(p_{t}^{V})\succeq-\beta I_{d}$ and $-\nabla^{2}\log(p_{t}^{2V})\succeq-\beta I_{d}$ uniformly over $\mathbb{R}^{d}$.
In order to control the gradient terms, we also need the converse bound across the whole trajectory. Fortunately, using the results of Mikulincer and Shenfeld, [36], smoothness at time $t=0$ suffices. We thus make the following additional assumption.
Assumption 2 (Smoothness)
The potential $V$ is $C^{2}(\mathbb{R}^{d})$ and has an $L$-Lipschitz gradient.
Proposition 9 (Mikulincer and Shenfeld, [36])
Under Assumption 2, it holds for all $z\in\mathbb{R}^{d}$ that
$\nabla^{2}\log(\pi)(z)-\nabla^{2}\log(p_{t}^{V})(z)\preceq\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}I_{d}\,.$ (5)
The proof is left to Appendix A.4 and is mainly an identification of terms. As a result, the following lower bound holds: $\nabla^{2}\log(p_{t}^{V})\succeq-(L+1)I_{d}\,.$ In particular, under Assumptions 1-2, we recover that $\|\nabla\log(p_{t}^{iV})(z)\|\leq(iL+\beta+1)\|z\|+\|\nabla\log(p_{t}^{iV})(0)\|$ for $i\in\{1,2\}$. Hence, in order to control the average quadratic error integrated w.r.t. $p_{t}^{V}$, it suffices to control the zeroth and the second order moments of the ratio $\Phi_{t}(z)\coloneqq\frac{p_{t}^{2V}}{p_{t}^{V}}$. This will be the object of the next paragraph.
3.2 Estimates on the moments of the ratio
In this subsection, we provide our upper bounds on the moments of $\Phi_{t}$. By slight abuse of notations, we shall denote $m_{i}(\Phi_{t})=\int\|x\|^{i}\Phi_{t}(z)\mathrm{d}z$ the $i$-th moment of $\Phi_{t}$.
Lemma 10 (Bound on the moments of the ratio)
Assume that $\mu\propto e^{-V}$ is $L$-log-smooth and $C^{2}$ and has finite second moment. It holds that
$\begin{cases}&m_{0}(\Phi_{t})\leq\frac{(Z_{V})^{2}}{Z_{2V}}e^{td}(L(1-e^{-2t})+e^{-2t})^{d}\,,\\ &m_{2}(\Phi_{t})\leq\frac{2e^{t(d+2)}(Z_{V})^{2}}{Z_{2V}}(L(1-e^{-2t})+e^{-2t})^{d+2}\left[m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\right]\,.\end{cases}$
The proof of Lemma 10 is provided in Appendix A.5 and relies on applying Grönwall’s lemma for the evolution of the moments. Crucially, note that this bound is only polynomial in $L$ when $d$ is fixed instead of exponential. This is a consequence of the bound of Mikulincer and Shenfeld, [36] that remains bounded with respect to $L$ whenever $t>0$. We have now all the ingredients to state our main result.
3.3 Assembling the pieces
We combine our non-asymptotic bound of Proposition 7 together with our previous estimates and Theorem 3 to recover a polynomial time bound in fixed dimension.
Theorem 11 (Main theorem)
Under Assumptions 1-2 and the additional assumptions that there exists $\xi\geq 0$ such that $\|\nabla\log(p_{t}^{V})(0)\|,\|\nabla\log(p_{t}^{2V})(0)\|\leq\xi$ and that $\mu$ has finite second moment $m_{2}(\mu)$, if we run algorithm Eq. 3 for an appropriate (explicit) set of parameters and with the stochastic score estimators $\hat{s}_{n_{k},t_{k}}$ defined in Eq. 4, then, denoting $\hat{p}$ the stochastic distribution of the output $Y_{N}$, it holds that
$\begin{cases}&\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon\,,\\ &\sum_{i=1}^{N}n_{k}\lesssim d^{7}L^{d+2}\epsilon^{-2(d+3)}(m_{2}(\mu)+d)^{2(d+1)}(K_{1}+K_{2})\,,\\ &K_{1}=(L+\beta+1)^{2}(L+1)^{2}\bigg{[}m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\bigg{]}\,,~{}~{}K_{2}=d\beta+\xi^{2}+1\,,\end{cases}$
where $\lesssim$ hides a universal constant as well as log factors with respect to $K_{1},K_{2}$.
The proof is left in Appendix A.9 and gives the appropriate set of parameters to use provided $L,R,m_{2}(\mu),\beta$ are known. We also provide in Appendix a version of this theorem that handles the case where the constants $L,R,\beta,m_{2}(\mu)$ are unknown.
Remark 12
As the proof shows, one could have relaxed the uniform-in-time upper-bound on $\nabla^{2}\log(p_{t}^{V})$, $\nabla^{2}\log(p_{t}^{2V})$ into a time-dependent version. This relaxation would simply have replaced the term $\beta$ in our bounds with $\max_{1\leq k\leq N}\beta_{t_{k}}$.
4 Application to smooth and strongly dissipative potentials
In this section, we prove that the results above apply in the strong dissipative case.
Assumption 3 (Strong Dissipativity)
The potential $V$ is $\alpha$-strongly convex outside some ball of radius $R$ centered at $m\in\mathbb{R}^{d}$ denoted $B_{R}(m)$.
We prove that if Assumptions 2-3 hold, then Assumption 1 holds with $\beta\asymp L^{2}(R^{2}(L/\alpha)^{2}+d/\alpha)$. The proof is decomposed into two time regimes: first, when $t$ is sufficiently small, the following result ensures that the forward process remains log-smooth, and a fortiori semi-log-concave, when $\mu$ is.
Proposition 13 (Chen et al., 2023a [10])
If $V$ is $C^{2}$ and $L$-smooth and if $t$ is such that $1-e^{-2t}\leq\frac{e^{-t}}{L}$, then it holds that $\|\nabla^{2}\log(p_{t}^{V})\|\leq 2Le^{t}$.
Hence, we only need to control $\nabla^{2}\log(p_{t}^{V})$ when $t\gtrsim 1$. Tweedie’s formula (see Appendix A.1) yields the identity $\nabla^{2}\log(p_{t}^{V})(x)-\nabla^{2}\log(\pi)(x)=\frac{e^{-2t}}{1-e^{-2t}}\left(\frac{\text{Cov}_{q_{t,x}}(X)}{1-e^{-2t}}-I_{d}\right)\,,$ where $q_{t,x}\propto e^{-V(z)}e^{-\frac{e^{-2t}\|z-xe^{t}\|^{2}}{2(1-e^{-2t})}}$. Hence, in order to control $\nabla^{2}\log(p_{t}^{V})$ from above, it suffices to upper-bound $\text{Cov}_{q_{t,x}}$ independently of $x$. Now remark that if $V$ is $\alpha$-strongly convex outside some ball of radius $R$ with condition number $\kappa\coloneqq L/\alpha$, then $q_{t,x}$ remains at least $\alpha$-strongly log-concave outside the same ball of radius $R$ and has most condition number $\kappa$. We thus derived the following result which upper-bounds the variance (and a fortiori the covariance matrix) of log-smooth and strongly log-concave measures outside some ball of radius $R$.
Proposition 14
Let $V$ be a potential that is $L$-smooth, $\alpha$-strongly convex outside some ball of radius $R$. Then, denoting $\mu\propto e^{-V}$ and $\kappa=L/\alpha$, it holds for any critical point $x^{*}$ that
$\text{Var}_{\mu}(X)\leq\mathbb{E}_{\mu}[\|X-x^{*}\|^{2}]\leq 8R^{2}\left(\kappa+1\right)^{2}+\frac{4d}{\alpha}\,.$
The proof is deferred to Appendix B.1.
Remark 15
We note that Ma et al., [31] and Chehab et al., [9] also derived bounds for the second moment of smooth and (weakly) dissipative potentials which a fortiori provides bounds for the variance. Yet the former suffers an exponential dependence in $LR^{2}$ and the latter implicitly depends on the the gradient of the potential at 00 which would have induced a dependence in $x$ when applied to $q_{t,x}$.
This upper-bound combined with Proposition 9 and Proposition 13 yields that the forward process $p_{t}^{V}$ is uniformly smooth along the trajectory.
Corollary 16
Assume $V$ is $C^{2}$, $L$-smooth and $\alpha$-strongly convex outside some ball of radius $R$ Then, denoting the condition number $\kappa=L/\alpha$, it holds that
$\|\nabla^{2}\log(p_{t}^{V})\|\leq 4(L+1)^{2}\left(8R^{2}\alpha(\kappa+1)^{2}+\frac{4d}{\alpha}+1\right)\,.$
The proof can be found in Appendix B.2. Note that numerous works in the reverse diffusion literature make the a priori assumption that the forward process is $L$-log-smooth [12, 24, 25]. The result above provides a non-trivial setting where this assumption does hold.
Remark 17
Note that our bounds extend the results of Mikulincer and Shenfeld, [36] and could be used in a similar fashion to recover estimates on the Log-Sobolev constants of strongly dissipative potentials. Yet, because of our residual dependence in $d$, we expect our resulting bounds to worsen those found in Ma et al., [31].
In order to apply our Theorem 11, it remains now to bound the second moment $m_{2}(\mu)$, the gradients of the forward process $\|\nabla\log(p_{t}^{V})\|,\|\nabla\log(p_{t}^{2V})\|$ and the density $\mu$ at 00 in the case where $V$ is $L$-smooth and $\alpha$-strongly convex outside some ball $B_{R}(m)$. In what follows, to ease notations, we shall assume that $\nabla V(0)=0$. Note that this extra assumption is benign since we can always compute any critical point $x^{*}$ up to arbitrary precision in polynomial time [23] and then sample from $\tilde{V}=V(\cdot+x^{*})$. Using Proposition 14, this extra assumption readily implies $m_{2}(\mu)\leq 8R^{2}\left(\kappa+1\right)^{2}+\frac{4d}{\alpha}$. We provide in Appendix B.3 polynomial bounds on the remaining quantities under Assumptions 2-3 which allows us to prove that strongly dissipative distributions can be sampled from in polynomial time in fixed dimension.
Corollary 18
Under Assumptions 2-3 and if we further assume $\nabla V(0)=0$, there exists a stochastic algorithm to sample from $\mu\propto e^{-V}$ that outputs a sample $Y\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon$ that runs in (deterministic) time $O(d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)})$ where $\lesssim$ hides a universal constant and log factors in $d,L,\alpha,R$.
The proof is left in Appendix B.4. We summarize in Table 1 how our bounds compare to those of previous works after plugging our estimate of $m_{2}(\mu)$ in the smooth, strongly dissipative case. Assuming that our estimate is tight, we observe that while our bound exhibits a worse dependence in $\epsilon^{-1},d$ than in Ma et al., [31] and in He et al., [22], we provide an exponential improvement with respect to $R,L$ over previous works where we lowered the complexity to polynomial. We provide in Appendix D a 2D simulation that compares standard ULA to our algorithm.
Algorithm Complexity
ULA [31] $O(e^{32LR^{2}}\kappa^{2}\frac{d}{\epsilon^{2}})$
RD + ULA [25] $e^{O(L^{3}\log^{3}(L^{2}(R^{2}\kappa^{2}+d/\alpha)/\epsilon))}$
RD + Rejection Sampling [22] $O(L^{d/2}\epsilon^{-d}e^{L\|x^{*}\|^{2}+L^{2}(\kappa^{2}R^{2}+d/\alpha)})$
RD + Self-normalized IS (ours) $O(d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)})$
Table 1: Complexity of sampling algorithms in the smooth, strongly dissipative case. We denoted $x^{*}$ the global minimum of $V$.
5 Conclusion
In this article, we pushed one step further the reduction from sampling to intermediate score estimation initiated over the past two years by Chen et al., 2023b [12], Chen et al., 2023a [10], Conforti et al., [16], Benton et al., [5]. Using our self-normalized estimator of the scores, we proved that in fixed dimension, the sampling problem could be further reduced to a regularity problem along the forward OU process. We then applied this result to derive polynomial time guarantees for the task of sampling from a log-smooth, strongly dissipative distribution thus providing an exponential improvement over previously known bounds. Yet we believe that our contribution could pave the way for numerous other non-log concave sampling problems in low dimensions. In future works, we plan to apply our results to Gaussian mixtures which are not fully covered by the strongly dissipative framework.
Acknowledgements
This work was funded by ANR JCJC WOS. We thank Pierre Monmarché for his help in proving Proposition 14 and Lemma 27.
Appendix A Proofs of Section 3
A.1 Identities of the forward process
Before starting our proofs, we recall useful identities on the density and the log hessian of the forward process known as the Tweedie’s formulas [39].
Proposition 19 (Tweedie’s formulas)
Denoting $p_{t}^{V}$ the density of the forward process $X_{t}$ initialized at $\mu\propto e^{-V}$, and $Y_{t}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$, it holds for all $z\in\mathbb{R}^{d}$ that
$p_{t}^{V}(z)=\frac{e^{td}}{Z^{V}}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]\,,$ (6)
and that
$\nabla^{2}\log(p_{t}^{V})(z)=\frac{\mathbb{E}[Y_{t}Y_{t}^{\top}e^{-V((z-Y_{t})e^{t})}]}{(1-e^{-2t})^{2}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]}-\frac{I_{d}}{(1-e^{-2t})}-(\nabla\log(p_{t}^{V})(z)))(\nabla\log(p_{t}^{V})(z))^{\top}\,.$ (7)
Proof.
Recall that $p_{t}^{V}$ is the law of the variable
$X_{t}=e^{-t}X_{0}+B_{1-e^{-2t}}\,,$
with $X_{0}\sim\mu$ and $B_{s}$ the standard Brownian motion evaluated at time $s$. Hence, using Bayes formula, we have
[…]
After taking the logarithm and differentiating with respect to $z$, we obtain
[…]
To obtain the Hessian, we differentiate the formula above. The Jacobian of the numerator is given by
[…]
from which we can deduce
[…]
$~{}~{}~{}-(\nabla\log(p_{t}^{V})(z))(\nabla\log(p_{t}^{V})(z))^{\top}\,.$
In order to rewrite the quantities above as expectations, we make the change of variable $y=z-xe^{-t}$ so that $x=(z-y)e^{t}$ and we obtain for the density $p_{t}^{V}$ :
$p_{t}^{V}(z)=\frac{e^{td}}{Z^{V}}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]\,,$
where $Y_{t}\sim\mathcal{N}(0,I_{d}(1-e^{-2t}))$. Conversely, the score rewrites as
$\nabla\log(p_{t}^{V})(z)=\frac{\mathbb{E}[-Y_{t}e^{-V((z-Y_{t})e^{t})}]}{(1-e^{-2t})\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]}\,,$
and the Hessian rewrites as
$\nabla^{2}\log(p_{t}^{V})(z)=\frac{\mathbb{E}[Y_{t}Y_{t}^{\top}e^{-V((z-Y_{t})e^{t})}]}{(1-e^{-2t})^{2}\mathbb{E}[e^{-V((z-Y_{t})e^{t})}]}-\frac{I_{d}}{(1-e^{-2t})}-(\nabla\log(p_{t}^{V})(z)))(\nabla\log(p_{t}^{V})(z))^{\top}\,.$
A.2 Proof of Proposition 6
Proof.
For the numerator, we have
$\hat{N}(z)-N(z)=\frac{-1}{n}\sum_{i=1}^{n}\frac{y_{i}e^{-V((z-y_{i})e^{t})}}{1-e^{-2t}}+N\,,$
hence, since the $y_{i},\;i=1,\dots,n$ are i.i.d. distributed as $Y_{t}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$,
$\mathbb{E}[\|\hat{N}(z)-N(z)\|^{2}]=\frac{1}{n}\mathbb{E}\left[\left\|\frac{Y_{t}e^{-V((z-Y_{t})e^{t})}}{1-e^{-2t}}-N(z)\right\|^{2}\right]\leq\frac{1}{n}\mathbb{E}\left[\frac{\|Y_{t}\|^{2}e^{-2V((z-Y_{t})e^{t})}}{(1-e^{-2t})^{2}}\right]\,.$
Taking in the trace in the log hessian identity in Proposition 19 yields
$\mathbb{E}\left[\frac{\|Y_{t}\|^{2}e^{-2V((z-Y_{t})e^{t})}}{(1-e^{-2t})^{2}}\right]=\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right)p_{t}^{2V}Z_{2V}e^{-td}\,.$
Similarly, we have
$\hat{D}(z)-D(z)=\frac{1}{n}\sum_{i=1}^{n}e^{-V((z-y_{i})e^{t})}-D\,,$
hence we get using again Proposition 19,
$\mathbb{E}[(\hat{D}(z)-D(z))^{2}]=\frac{1}{n}\left(\mathbb{E}[e^{-2V((z-Y_{t})e^{t})}]-\mathbb{E}[e^{-2((z-Y_{t})e^{t})}]^{2}\right)\leq\frac{p_{t}^{2V}(z)Z_{2V}e^{-td}}{n}\,.$ (8)
A.3 Proof of Proposition 7
Proof.
To ease notations, we shall drop the dependence in $z$ throughout the proof. Define the event $A=(\hat{D}\geq\eta D)\cap(\|\hat{N}\|\leq\kappa\|N\|)$ where $\eta\leq 1,\kappa\geq 1$ are positive scalars to be chosen later. We start to decompose the quadratic error as:
$\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\right]=\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{A}\right]+\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{\bar{A}}\right]\,.$
We now separately analyze the first and the second term. For the first term, define
$\begin{cases}\theta:\mathbb{R}^{d}\times\mathbb{R}^{*}\to\mathbb{R}\\ (x,p)\mapsto\left\|\frac{x}{p}-\frac{N}{D}\right\|^{2}\,.\end{cases}$
The gradient and Hessian of $\theta$ are given by
$\begin{cases}&\nabla\theta(x,p)=\frac{-2}{p}\left(\frac{N}{D}-\frac{x}{p},\left\|\frac{x}{p}\right\|^{2}-\left\langle\frac{x}{p},\frac{N}{D}\right\rangle\right)\,,\\ &\nabla^{2}\theta(x,p)=\frac{-2}{p^{2}}\begin{pmatrix}-I_{d}&(\frac{2x}{p}-\frac{N}{D})^{\top}\\ (\frac{2x}{p}-\frac{N}{D})&-3\left\|\frac{x}{p}\right\|^{2}+2\langle\frac{x}{p},\frac{N}{D}\rangle\end{pmatrix}\,.\end{cases}$
We thus make a Taylor expansion of order $2$ of $\theta(\hat{N},\hat{D})$ around $(N,D)$: there exists (a random) $\hat{t}\in[0,1]$ such that
$\begin{split}\theta(\hat{N},\hat{D})&=\theta(N,D)+\nabla\theta(N,D)^{\top}(\hat{N}-N,\hat{D}-D)\\ &~{}~{}~{}+\frac{1}{2}(\hat{N}-N,\hat{D}-D)^{\top}\nabla^{2}\theta(\hat{N}_{\hat{t}},\hat{D}_{\hat{t}})(\hat{N}-N,\hat{D}-D)\,.\end{split}$
where we denoted $\hat{N}_{\hat{t}}=\hat{t}\hat{N}+(1-\hat{t})N$ and $\hat{D}_{\hat{t}}=\hat{t}\hat{D}+(1-\hat{t})D$. The two first terms in the expansion are null and we are left with
$\theta(\hat{N},\hat{D})$ $=\frac{1}{\hat{D}_{\hat{t}}^{2}}\bigg{(}\|\hat{N}-N\|^{2}-2(\langle 2\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}}-\frac{N}{D},\hat{N}-N\rangle(\hat{D}-D))+\left(3\frac{\|\hat{N}_{\hat{t}}\|^{2}}{\hat{D}_{\hat{t}}^{2}}-2\langle\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}},\frac{N}{D}\rangle\right)(\hat{D}-D)^{2}\bigg{)}$
$\leq\frac{1}{\hat{D}_{\hat{t}}^{2}}\left(\|\hat{N}-N\|^{2}+2\|2\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}}-\frac{N}{D}\|\|\hat{N}-N\||\hat{D}-D|+\left(3\frac{\|\hat{N}_{\hat{t}}\|^{2}}{\hat{D}_{\hat{t}}^{2}}+2\left\|\frac{\hat{N}_{\hat{t}}}{\hat{D}_{\hat{t}}}\right\|\left\|\frac{N}{D}\right\|\right)(\hat{D}-D)^{2}\right)\,.$
Hence, almost surely over $A$
$\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\leq\frac{1}{\eta^{2}D^{2}}\left(\|\hat{N}-N\|^{2}+\frac{6\kappa}{\eta}\left\|\frac{N}{D}\right\|\|\hat{N}-N\||\hat{D}-D|+5\left\|\frac{N}{D}\right\|^{2}\left(\frac{\kappa}{\eta}\right)^{2}(\hat{D}-D)^{2}\right)\,.$
Hence, after taking the expectation and applying Cauchy-Schwarz, we obtain
$\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{A}\right]\leq\frac{1}{\eta^{2}D^{2}}$ $(\mathbb{E}[\|\hat{N}-N\|^{2}]+\frac{6\kappa}{\eta}\left\|\frac{N}{D}\right\|\mathbb{E}[\|\hat{N}-N\|^{2}]^{1/2}\mathbb{E}[(\hat{D}-D)^{2}]^{1/2}$
$~{}~{}~{}+5\left\|\frac{N}{D}\right\|^{2}\left(\frac{\kappa}{\eta}\right)^{2}\mathbb{E}[(\hat{D}-D)^{2}])\,.$
Now recall that $D=p_{t}^{V}Z_{V}e^{-td}$ and that $\frac{N}{D}=\nabla\log(p_{t}^{V})$ which, combined with Proposition 6 yields for the first term:
$\frac{\mathbb{E}[\|\hat{N}-N\|^{2}]}{\eta^{2}D^{2}}\leq\frac{p_{t}^{2V}Z_{2V}e^{td}}{\eta^{2}(p_{t}^{V})^{2}(Z_{V})^{2}n}\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right)\,,$
for the second term:
$\frac{6\kappa}{\eta^{3}D^{2}}\left\|\frac{N}{D}\right\|\|\hat{N}-N\||\hat{D}-D|\leq\frac{6\kappa p_{t}^{2V}Z_{2V}e^{td}}{\eta^{3}(p_{t}^{V})^{2}(Z_{V})^{2}n}\|\nabla\log(p_{t}^{V})\|\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right)^{1/2}$
and for the last term:
$5\left\|\frac{N}{D}\right\|^{2}\left(\frac{\kappa}{\eta^{2}D}\right)^{2}(\hat{D}-D)^{2}\leq\frac{5\kappa^{2}p_{t}^{2V}Z_{2V}e^{td}}{\eta^{4}(p_{t}^{V})^{2}(Z_{V})^{2}n}\|\nabla\log(p_{t}^{V})\|^{2}\,.$
Hence we finally obtain
$\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{A}\right]\leq\frac{p_{t}^{2V}Z_{2V}e^{td}}{n\eta^{2}(p_{t}^{V}Z_{V})^{2}}$ $\bigg{(}\frac{6\kappa}{\eta}\|\nabla\log(p_{t}^{V})\|\left[\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right]^{1/2}$
$~{}~{}~{}+\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+6\left(\frac{\kappa}{\eta}\right)^{2}\|\nabla\log(p_{t}^{V})\|^{2}\bigg{)}\,.$ (9)
Let us now handle the quadratic error of the estimator on the complementary $\bar{A}$. We have, using Young’s inequality $||a-b||^{2}\leq 2(||a||^{2}+||b||^{2})$,
$\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{\bar{A}}\right]\leq 2\left\|\frac{N}{D}\right\|^{2}\mathbb{P}(\bar{A})+2\mathbb{E}\left[\mathds{1}_{\bar{A}}\frac{\max_{i}\|y_{i}\|^{2}}{(1-e^{-2t})^{2}}\right]\,.$
Now, recall that $X_{i}=||y_{i}||^{2}(1-e^{-2t})^{-1}$ are $n$ independent variables such that for all $i$, $X_{i}\sim\chi^{2}(d)$. Using Hölder inequality for some $p\geq 1$, the second term can be upper-bounded as
$\mathbb{E}\left[\mathds{1}_{\bar{A}}\frac{\max_{i}\|y_{i}\|^{2}}{(1-e^{-2t})^{2}}\right]$ $\leq\frac{1}{1-e^{-2t}}\mathbb{E}[\max_{i}X_{i}^{p}]^{1/p}\mathbb{P}(\bar{A})^{1-1/p}$
$\leq\frac{1}{1-e^{-2t}}(n\mathbb{E}[X_{1}^{p}])^{1/p}\mathbb{P}(\bar{A})^{1-1/p}\leq\frac{1}{1-e^{-2t}}n^{1/p}(d+2p)\mathbb{P}(\bar{A})^{1-1/p}\,,$
where we used in the penultimate inequality that the max is smaller than the sum, and in the last one that $\mathbb{E}[X_{1}^{p}]=\prod_{i=0}^{p-1}(d+2i)$ when $X_{1}\sim\chi^{2}(d)$ combined with the fact that the geometric mean is lower than the arithmetic mean.
We now upper bound the probability of the event $\bar{A}=(\hat{D}<\eta D)\cup(\|\hat{N}\|>\kappa\|N\|)$. By Tchebychev inequality, using $\eta<1$, it holds that
$\mathbb{P}(\hat{D}<\eta D)$ $\leq\frac{\mathbb{E}[(\hat{D}-D)^{2}]}{D^{2}(\eta-1)^{2}}\leq\frac{p_{t}^{2V}Z_{2V}e^{td}}{n(p_{t}^{V}Z_{V})^{2}(\eta-1)^{2}}\coloneqq\frac{U}{n(\eta-1)^{2}}\,.$
Similarly, recalling that $||N||=D||\nabla\log(p_{t}^{V})||$, we have
$\mathbb{P}(\|\hat{N}\|>\kappa\|N\|)$ $\leq\frac{\mathbb{E}[||\hat{N}-N||^{2}]}{||N||^{2}(\kappa-1)^{2}}$
$\leq\frac{U}{n\|\nabla\log(p_{t}^{V})\|^{2}(\kappa-1)^{2}}\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}\right)\,.$
We now make a disjunction of cases: if $\|\nabla\log(p_{t}^{V})\|\geq 1$, we pick $\eta=1/2$ and $\kappa=3/2$ so we recover
$\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{\bar{A}}\right]$ $\leq\frac{8U}{n}\left[\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+2\|\nabla\log(p_{t}^{V})\|^{2}\right]$
$~{}~{}+\frac{n^{1/p}(d+2p)}{1-e^{-2t}}\left[\frac{8U}{n}\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+1\right)\right]^{1-1/p}$
$\leq\frac{8U}{n}\left[\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+2\|\nabla\log(p_{t}^{V})\|^{2}\right]$
$~{}~{}+\frac{n^{1/p}(d+2p)}{1-e^{-2t}}\left[\frac{8U}{n}\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+1\right)+1\right]\,.$
We thus pick $p=\log(n)$ to get
$\begin{split}\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{\bar{A}}\right]\leq\frac{16e^{2}(d+2\log(n))}{n(1-e^{-2t})}\left[U\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+1\right)+1\right]\,.\end{split}$
Combining this with the bound Eq. A.3 eventually yields
$\delta^{2}\leq\frac{32e^{2}(d+2\log(n))}{n(1-e^{-2t})}\left[U\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+\|\nabla\log(p_{t}^{2V})\|^{2}+1\right)+1\right]\,,$
where we used the inequality
$~{}~{}~{}~{}~{}\|\nabla\log(p_{t}^{V})\|\left[\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right]^{1/2}$
$\leq\frac{1}{2}\left(\|\nabla\log(p_{t}^{V})\|^{2}+\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right)\,.$
In the case where $\|\nabla\log(p_{t})\|^{2}<1$, we instead pick $\eta=1/2$ and $\kappa=1+\frac{1}{2\|\nabla\log(p_{t}^{V})\|}$. We obtain that
$\mathbb{P}(\bar{A})\leq\frac{4U}{n}\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+1\right)$
and as previously, for $p=\log(n)$ we get
$\begin{split}\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{\bar{A}}\right]\leq\frac{16e^{2}(d+2\log(n))}{n(1-e^{-2t})}\left[U\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+1\right)+1\right]\,.\end{split}$
For this choice of $\kappa,\eta$, the bound on $A$ becomes
$\mathbb{E}\left[\left\|\frac{\hat{N}}{\hat{D}}-\frac{N}{D}\right\|^{2}\mathds{1}_{A}\right]$ $\leq\frac{4U}{n}\bigg{(}6\left[\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{2V})\|^{2}\right]^{1/2}+\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+6\bigg{)}$
$\leq\frac{4U}{n}\left(7(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}})+6(\|\nabla\log(p_{t}^{2V})\|^{2}+1)\right)\,.$
Thus, we obtain as previously
$\delta^{2}\leq\frac{32e^{2}(d+2\log(n))}{n(1-e^{-2t})}\left[U\left(\Delta\log(p_{t}^{2V})-\frac{\Delta\log(\pi)}{1-e^{-2t}}+\|\nabla\log(p_{t}^{V})\|^{2}+\|\nabla\log(p_{t}^{2V})\|^{2}+1\right)+1\right]\,,$
A.4 Proof of Proposition 9
Proof.
Define the Ornstein-Uhlenbeck semi-group $Q_{t}$ as
$Q_{t}(g)(z)$ $=\int g(ze^{-t}+\sqrt{1-e^{-2t}}y)e^{-\frac{\|y\|^{2}}{2}}(2\pi)^{-d/2}\mathrm{d}y$
$=\frac{1}{\sqrt{1-e^{-2t}}}\int g(u)e^{-\frac{\|u-e^{-t}z\|^{2}}{2(1-e^{-2t})}}(2\pi)^{-d/2}\mathrm{d}u\,,$
for all function $g$ integrable w.r.t. the standard Gaussian measure. Taking $g$ as $f=\frac{\mathrm{d}\mu}{\mathrm{d}\pi}$ with $\pi$ the standard Gaussian, we obtain that
$Q_{t}(f)(z)$ $=\frac{1}{Z_{V}\sqrt{1-e^{-2t}}}\int e^{-V(u)}e^{\frac{\|u\|^{2}}{2}}e^{-\frac{\|u-e^{-t}z\|^{2}}{2(1-e^{-2t})}}\mathrm{d}u$
$=\frac{1}{Z_{V}\sqrt{1-e^{-2t}}}\int e^{-V(u)}e^{\frac{\|u\|^{2}(1-e^{-2t})-\|u\|^{2}+\langle z,ue^{-t}\rangle-e^{-2t}\|z\|^{2}}{2(1-e^{-2t})}}\mathrm{d}u$
$=\frac{1}{Z_{V}\sqrt{1-e^{-2t}}}\int e^{-V(u)}e^{\frac{-\|ue^{-t}-z\|^{2}+\|z\|^{2}-e^{-2t}\|z\|^{2}}{2(1-e^{-2t})}}\mathrm{d}u$
$=\frac{e^{\frac{\|z\|^{2}}{2}}}{Z_{V}\sqrt{1-e^{-2t}}}\int e^{-V(u)}e^{-\frac{\|ue^{-t}-z\|^{2}}{2(1-e^{-2t})}}\mathrm{d}u\,.$
In particular, we remark that $\nabla\log(Q_{t}(f))=\nabla\log(p_{t}^{V})-\nabla\log(\pi)$. Now, the quantity $\nabla\log(Q_{t}(f))$ was studied in Mikulincer and Shenfeld, [35] and they prove in Lemma 5 that for all $z$
$\nabla^{2}\log(Q_{t}(f))(z)\succeq\frac{(1-L)e^{-2t}}{(1-e^{-2t})(L-1)+1}I_{d}\,,$
which is equivalent to
$\nabla^{2}\log(\pi)(z)-\nabla^{2}\log(p_{t}^{V})(z)\preceq\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}I_{d}\,.$
A.5 Proof of Lemma 10
Before proving Lemma 10, we introduce this preliminary result on the evolution of $\Phi_{t}$.
Lemma 20 (Evolution of the ratio)
Let $t>0$, it holds that
$\partial\Phi_{t}=\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\,.$
Proof.
Recall that the log-density $\log(p_{t}^{V})$ evolves as
$\partial\log(p_{t}^{V})=\Delta\log(p_{t}^{V})+\|\nabla\log(p_{t}^{V})\|^{2}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)\,.$
Hence, we deduce that $\log(\Phi_{t})$ evolves as
$\partial\log(\Phi_{t})=\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(p_{t}^{2V})\|^{2}-\|\nabla\log(p_{t}^{V})\|^{2}\,.$
The difference of quadratic terms can be expressed as
$\|\nabla\log(p_{t}^{2V})\|^{2}-\|\nabla\log(p_{t}^{V})\|^{2}$ $=\|\nabla\log(\Phi_{t})+\nabla\log(p_{t}^{V})\|^{2}-\|\nabla\log(p_{t}^{V})\|^{2}$
$=\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\,,$
which allows to recover
$\partial\log(\Phi_{t})=\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\,.$
We now provide the proof of Lemma 10.
Proof.
Until the rest of the proof, the dependence on $z$ of the integrand shall be implied unless expressed explicitly. We start by differentiating $m_{0}(\Phi_{t})$ with respect to $t$:
$\partial m_{0}(\Phi_{t})$ $=\int\partial\Phi_{t}\mathrm{d}z$
$=\int\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\mathrm{d}z\,,$
where we used Lemma 20 to compute $\partial\Phi_{t}$. Using integration by parts, the first term reads
$\int\Phi_{t}\Delta\log(\Phi_{t})\mathrm{d}z$ $=-\int\langle\nabla\Phi_{t},\nabla\log(\Phi_{t})\rangle\mathrm{d}z=-\int\Phi_{t}\|\nabla\log(\Phi_{t})\|^{2}\mathrm{d}z,$
hence the first and the third terms cancel and we recover
$\partial m_{0}(\Phi_{t})=2\int\langle\nabla\log(p_{t}^{V}),\nabla\Phi_{t}\rangle\mathrm{d}z-\int\langle\nabla\log(\pi),\nabla\Phi_{t}\rangle\mathrm{d}z\,.$
Using integration by parts again, we recover
$\partial m_{0}(\Phi_{t})$ $=\int\Delta\log(\pi)\Phi_{t}\mathrm{d}z-2\int\Delta\log(p_{t}^{V})\Phi_{t}\mathrm{d}z$
$=2\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\Phi_{t}\mathrm{d}z-\int\Delta\log(\pi)\Phi_{t}\mathrm{d}z$
$=dm_{0}(\Phi_{t})+2\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\Phi_{t}\mathrm{d}z\,.$
Using Proposition 9, since $V$ is $L$-smooth, the term $(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))$ can be upper-bounded uniformly by $\frac{d(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}$ so we eventually get
$\partial m_{0}(\Phi_{t})\leq dm_{0}(\Phi_{t})\left(1+\frac{2(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}\right)\,.$
Hence we can use Gronwall lemma which yields
$m_{0}(\Phi_{t})\leq m_{0}(\Phi_{0})\exp\left(d\int_{0}^{t}\left(1+\frac{2(L-1)e^{-2s}}{(1-e^{-2s})(L-1)+1}\right)\mathrm{d}s\right)\,.$
Denoting by $Z_{V}$ (resp. $Z_{2V}$) the normalizing constant of $e^{-V}$ (resp. $e^{-2V}$), the term $m_{0}(\Phi_{0})$ reads
$m_{0}(\Phi_{0})$ $=\int\frac{p_{0}^{2V}(z)}{p_{0}^{V}(z)}\mathrm{d}z=\frac{Z_{V}}{Z_{2V}}\int\frac{e^{-2V(z)}}{e^{-V(z)}}\mathrm{d}z=\frac{(Z_{V})^{2}}{Z_{2V}}\,.$
Finally, let us compute the integral above. Making the change of variable $u=e^{-2s}(L-1)$ we have $\mathrm{d}u=-2(L-1)e^{-2s}\mathrm{d}s$ which yields
$\int_{0}^{t}\frac{2(L-1)e^{-2s}}{(1-e^{-2s})(L-1)+1}\mathrm{d}s$ $=-\int_{L-1}^{(L-1)e^{-2t}}\frac{1}{L-u}\mathrm{d}u$
$=\left[\log(L-u)\right]_{L-1}^{(L-1)e^{-2t}}$
$=\log(L-(L-1)e^{-2t})$
$=\log(L(1-e^{-2t})+e^{-2t})\,.$
Hence we recover
$m_{0}(\Phi_{t})\leq\frac{e^{td}(Z_{V})^{2}}{Z_{2V}}(L(1-e^{-2t})+e^{-2t})^{d}\,.$
A.6 Upper bound on the second moment of the ratio
In order to recover a bound on the second moment of $\Phi_{t}$, we need several intermediate results. We first prove that the maximum of the ratio decreases through time.
Lemma 21 (Decrease of the maximum of $\Phi$)
The maximum of the ratio $\Phi_{t}$ decreases with $t$.
Proof.
Let $z_{t}$ be a point where $\Phi_{t}$ attains its maximum and denote $M_{t}=\log(\Phi_{t})(z_{t})$. By the implicit function theorem, $z_{t}$ is differentiable hence we can compute $\partial M_{t}$ as
$\partial M_{t}$ $=\partial\log(\Phi_{t})(z_{t})+\langle\partial z_{t},\nabla\log(\Phi_{t})(z_{t})\rangle$
$=\Delta\log(\Phi_{t})(z_{t})\,.$
Since $z_{t}$ is a maximum, we have in particular $\Delta\Phi_{t}(z_{t})\leq 0$ which implies that $M_{t}$ decreases. ∎
We then derive an upper-bound on the maximum of a log-smooth distribution.
Proposition 22 (Upper-bound of the maximum)
Let $V$ be a $C^{2}$ potential such that $\nabla^{2}V\preceq LI_{d}$. Denoting $\mu\propto e^{-V}$, it holds that $\frac{\mathrm{d}\mu}{\mathrm{d}z}\leq\left(\frac{\beta}{2\pi}\right)^{\frac{d}{2}}$.
Proof.
Recall that the density of $\mu$ can be re-written as
$\frac{\mathrm{d}\mu}{\mathrm{d}z}=\frac{e^{-(V(z)-V_{*})}}{\int_{z}e^{-(V(z)-V_{*})}\mathrm{d}z}\,,$
where $V_{*}$ the minimum of $V$ attained for some $z_{*}$. By definition $e^{-(V(z)-V_{*})}\leq 1$ for all $z$. Furthermore, since $V$ verifies $\nabla^{2}V\preceq LI_{d}$, we are ensured that
$V(z)-V_{*}\leq L\frac{\|z-z_{*}\|^{2}}{2}\,,$
which implies in particular that
$\frac{1}{\int_{z}e^{-(V(z)-V_{*})}\mathrm{d}z}\leq\left(\frac{L}{2\pi}\right)^{\frac{d}{2}}\,.$
Using the previous result, we can derive an upper-bound on the integrated squared gradient at 00.
Lemma 23 (Upper-bound integrated gradient)
Let $\mu\propto e^{-V}$ be a $L$-log smooth measure. Denoting $\mu(0)$ the density of $\mu$ with respect to the Lebesgue measure at 00, it holds that
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s\leq-\log\left(\mu(0)\right)+\frac{d}{2}\log\left(\frac{L}{2\pi}\right)\,.$
Proof.
Denoting $\pi$ the density of the standard $d$ dimensional Gaussian, recall that the density $p_{t}^{V}$ evolves as
$\partial p_{t}^{V}=\nabla\cdot\left(p_{t}^{V}\nabla\log\left(\frac{p_{t}^{V}}{\pi}\right)\right)\,,$
which can also be re-written as
$\partial\log(p_{t}^{V})$ $=\frac{\Delta p_{t}^{V}}{p_{t}^{V}}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)$
$=\Delta\log(p_{t}^{V})+\|\nabla\log(p_{t}^{V})\|^{2}-\langle\nabla\log(p_{t}^{V}),\nabla\log(\pi)\rangle-\Delta\log(\pi)\,.$
In particular, for $z=0$ this yields
$\partial\log(p_{t}^{V})(0)=\Delta\log(p_{t}^{V})(0)+\|\nabla\log(p_{t}^{V})\|^{2}(0)-\Delta\log(\pi)(0)\,,$
which implies
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s=\log(p_{t}^{V})(0)-\log(p_{0}^{V})(0)+\int_{0}^{t}\Delta\log(\pi)(0)-\Delta\log(p_{t}^{V})(0)\mathrm{d}s\,.$
Using the uniform upper-bound of Proposition 9, the second term can be upper-bounded as
$\int_{0}^{t}\Delta\log(\pi)(0)-\Delta\log(p_{t}^{V})(0)\mathrm{d}s$ $\leq d\int_{0}^{t}\frac{(L-1)e^{-2s}}{(1-e^{-2s})(L-1)+1}\mathrm{d}s$
$=\frac{d}{2}(L-1)\int_{e^{-2t}}^{1}\frac{1}{L-u(L-1)}\mathrm{d}u$
$=\frac{d}{2}\log(L(1-e^{-2t})+e^{-2t})\,.$
Furthermore, Proposition 9 shows that $\nabla^{2}\log(p_{t}^{V})\preceq\frac{L}{L(1-e^{-2t})+e^{-2t}}$. Thus, using Proposition 22, we recover that $\log(p_{t}^{V})(0)\leq\frac{d}{2}\log\left(\frac{L}{L(1-e^{-2t})+e^{-2t}}\right)-\frac{d}{2}\log(2\pi)$. In particular, we recover
$\int_{0}^{t}\|\nabla\log(p_{s}^{V})\|^{2}(0)\mathrm{d}s\leq-\log\left(\frac{\mathrm{d}\mu}{\mathrm{d}x}(0)\right)+\frac{d}{2}\log(L)-\frac{d}{2}\log(2\pi)\,.$
We can now bound the first order moment of $\Phi_{t}$
Lemma 24
Let $V$ be a $C^{2}$, $L$-smooth potential such that $\mu\propto e^{-V}$ has finite second moment $m_{2}(\mu)$. It holds that
$m_{1}(\Phi_{t})\leq\frac{(Z_{V})^{2}}{Z_{2V}}e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d}\left(\sqrt{m_{2}(\mu)}+\sqrt{-2\log\left(\mu(0)\right)+d\log\left(\frac{L}{2\pi}\right)}+2\sqrt{dL}\right)\,.$
Proof.
We differentiate $m_{1}(\Phi_{t})$ and we recover
$\partial m_{1}(\Phi_{t})=\int\Phi_{t}\left(\Delta\log(\Phi_{t})-\langle\nabla\log(\Phi_{t}),\nabla\log(\pi)\rangle+\|\nabla\log(\Phi_{t})\|^{2}+2\langle\nabla\log(p_{t}^{V}),\nabla\log(\Phi_{t})\rangle\right)\|z\|\mathrm{d}z\,.$
Integration by parts of the first term yields
$\int\Phi_{t}\Delta\log(\Phi_{t})\|z\|\mathrm{d}z=-\int\Phi_{t}\|\nabla\log(\Phi_{t})\|^{2}\|z\|+\langle\nabla\Phi_{t},\frac{z}{\|z\|}\rangle\mathrm{d}z\,,$
hence the squared gradients terms cancel and we recover
$\partial m_{1}(\Phi_{t})=\int\langle 2\nabla\log(p_{t}^{V})-\nabla\log(\pi),\nabla\Phi_{t}\rangle\|z\|\mathrm{d}z-\int\langle\nabla\Phi_{t},\frac{z}{\|z\|}\rangle\mathrm{d}z\,.$
Let us denote by $A$ the first term above. Integration by parts yields:
$A$ $=\int\Phi_{t}(\Delta\log(\pi)-2\Delta\log(p_{t}^{V}))\|z\|\mathrm{d}z+\int\Phi_{t}\langle\nabla\log(\pi)-2\nabla\log(p_{t}^{V}),\frac{z}{\|z\|}\rangle\mathrm{d}z$
$=-\int\Phi_{t}\Delta\log(\pi)\|z\|\mathrm{d}z+2\int\Phi_{t}(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|\mathrm{d}z$
$~{}~{}~{}-\int\Phi_{t}\langle\nabla\log(\pi),\frac{z}{\|z\|}\rangle\mathrm{d}z+2\int\Phi_{t}\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),\frac{z}{\|z\|}\rangle\mathrm{d}z$
$=(d+1)m_{1}(\Phi_{t})+2\int\Phi_{t}(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|\mathrm{d}z+2\int\Phi_{t}\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),\frac{z}{\|z\|}\rangle\mathrm{d}z\,.$
Using the upper bound given in Proposition 9, we get $2\int\Phi_{t}(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|\mathrm{d}z\leq 2d\frac{(L-1)e^{-2t}}{(L-1)(1-e^{-2t})+1}m_{1}(\Phi_{t})$. Similarly, we re-write the second term as
$2\int\Phi_{t}\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),\frac{z}{\|z\|}\rangle\mathrm{d}z$ $=2\int\Phi_{t}\langle\nabla\log\left(\frac{\pi}{p_{t}^{V}}\right)(z)-\nabla\log\left(\frac{\pi}{p_{t}^{V}}\right)(0),\frac{z}{\|z\|}\rangle\mathrm{d}z$
$~{}~{}~{}+2\int\Phi_{t}\langle\nabla\log\left(\frac{\pi}{p_{t}^{V}}\right)(0),\frac{z}{\|z\|}\rangle\mathrm{d}z$
$\leq 2m_{1}(\Phi_{t})\frac{(L-1)e^{-2t}}{(L-1)(1-e^{-2t})+1}+2\|\nabla\log(p_{t}^{V})(0)\|m_{0}(\Phi_{t})\,.$
Let us now handle the term $B=-\int\langle\nabla\Phi_{t},\frac{z}{\|z\|}\rangle\mathrm{d}z$. In one dimension, $B=2\Phi_{t}(0)\leq\max(\Phi_{t})$ and for $d\geq 2$, we have
$B$ $=\int\frac{\Phi_{t}(d-1)}{\|z\|}\mathrm{d}z$
$=\int_{B_{R}}\frac{\Phi_{t}(d-1)}{\|z\|}\mathrm{d}z+\int_{\overline{B_{R}}}\frac{\Phi_{t}(d-1)}{\|z\|}\mathrm{d}z$
$\leq\max(\Phi_{t})\int_{B_{R}}\frac{d-1}{\|z\|}\mathrm{d}z+\frac{d-1}{R}m_{0}(\Phi_{t})$
$=\max(\Phi_{t})\frac{2\pi^{d/2}}{\Gamma(d/2)}R^{d-1}+\frac{d-1}{R}m_{0}(\Phi_{t})\,.$
Using Lemma 21 and Proposition 22, we have that $\max(\Phi_{t})\leq\max(\Phi_{0})=\frac{(Z_{V})^{2}}{Z_{2V}}\max(\frac{\mathrm{d}\mu}{\mathrm{d}x})\leq\frac{(Z_{V})^{2}}{Z_{2V}}(\frac{L}{2\pi})^{d/2}$. Hence, if we pick $R=\left(Z_{2V}m_{0}(\Phi_{t})\Gamma(d/2)/Z_{V}^{2}\right)^{1/d}L^{-1/2}2^{1/2-1/d}$, we get as an upper-bound for $B$:
$B$ $\leq\frac{Z_{V}^{2}}{Z_{2V}}d\sqrt{L}2^{1/d-1/2}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\Gamma(d/2)^{-1/d}$
$\leq\frac{2Z_{V}^{2}}{Z_{2V}}\sqrt{dL}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\,.$
In particular, we recover that
$\partial m_{1}(\Phi_{t})\leq(d+1)\left(1+2\frac{(L-1)e^{-2t}}{(L-1)(1-e^{-2t})+1}\right)m_{1}(\Phi_{t})+2\|\nabla\log(p_{t}^{V})(0)\|m_{0}(\Phi_{t})+\frac{2Z_{V}^{2}}{Z_{2V}}\sqrt{Ld}(m_{0}(\Phi_{t})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\,,$
hence using Gronwall lemma, we have that
$m_{1}(\Phi_{t})$ $\leq e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}m_{1}(\Phi_{0})$
$~{}~{}~{}+2\int_{0}^{t}\left[\|\nabla\log(p_{s}^{V})(0)\|m_{1}(\Phi_{s})+\sqrt{dL}(m_{0}(\Phi_{s})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\right]\frac{e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}}{e^{s(d+1)}(L(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s\,.$
Using Lemma 10, we have that $m_{0}(\Phi_{s})\leq\frac{Z_{V}^{2}}{Z_{2V}}e^{sd}(L(1-e^{-2s})+e^{-2s})^{d}$ hence the first term of the integral is upper-bounded as:
$\int_{0}^{t}\frac{e^{-s(d+1)}m_{0}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|}{(L(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s\leq\frac{Z_{V}^{2}}{Z_{2V}}\int_{0}^{t}\|\nabla\log(p_{s}^{V})(0)\|\frac{e^{-s}}{L(1-e^{-2s})+e^{-2s}}\mathrm{d}s\,.$
By Cauchy-Schwarz it holds that
$\int_{0}^{t}\frac{\|\nabla\log(p_{s}^{V})(0)\|e^{-s}}{L(1-e^{-2s})+e^{-2s}}\mathrm{d}s\leq\sqrt{\int_{0}^{t}\|\nabla\log(p_{s}^{V})(0)\|^{2}\mathrm{d}s}\sqrt{\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s}\,.$
The integral term is given by
$\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s$ $=\frac{1}{2}\int_{e^{-2t}}^{1}\frac{1}{(L(1-u)+u)^{2}}\mathrm{d}u$
$=\frac{1}{2(1-L)}[-\frac{1}{L(1-u)+u}]_{e^{-2t}}^{1}$
$=\frac{1}{2(1-L)}(\frac{1}{L(1-e^{-2t})+e^{-2t}}-1)$
$=\frac{1-e^{-2t}}{2}$
Similarly,
$\int_{0}^{t}m_{0}(\Phi_{s})^{\frac{d-1}{d}}\frac{e^{-s(d+1)}}{(L(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s$ $\leq\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s$
$=\frac{1-e^{-2t}}{2}$
Hence, we obtain
$m_{1}(\Phi_{t})\leq\frac{Z_{V}^{2}}{Z_{2V}}e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}\left(\frac{Z_{2V}}{Z_{V}^{2}}m_{1}(\Phi_{0})+\sqrt{-2\log(\mu(0))+d\log\left(\frac{L}{2\pi}\right)}+\sqrt{dL}\right)\,.$
Finally, $m_{1}(\Phi_{0})=\frac{(Z_{V})^{2}}{Z_{2V}}\int\|z\|\frac{e^{-V(z)}}{\int e^{-V}\mathrm{d}z}\mathrm{d}z=\frac{(Z_{V})^{2}}{Z_{2V}}\sqrt{m_{2}(\mu)}$. ∎
We can derive our upper-bound on $m_{2}(\Phi_{t})$.
Proof.
We start by differentiating $m_{2}(\Phi_{t})$:
$\partial m_{2}(\Phi_{t})$ $=\int\|z\|^{2}\partial\Phi_{t}\mathrm{d}z$
$=\int\|z\|^{2}(\text{div}(\nabla\Phi_{t})-\langle\nabla\Phi_{t},\nabla\log(\pi)\rangle+2\langle\nabla\log(p_{t}^{V}),\nabla\Phi_{t}\rangle)\mathrm{d}z$
$=-\int 2\langle z,\nabla\Phi_{t}\rangle\mathrm{d}z+\int(\Delta\log(\pi)-2\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z+2\int\langle\nabla\log(\pi)-2\log(p_{t}^{V}),z\rangle\Phi_{t}\mathrm{d}z$
$=-2\int\Delta\log(\pi)\Phi_{t}\mathrm{d}z-\int\Delta\log(\pi)\|z\|^{2}\Phi_{t}\mathrm{d}z-2\int\langle\nabla\log(\pi),z\rangle\Phi_{t}\mathrm{d}z$
$~{}~{}~{}+2\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z+4\int\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),z\rangle\Phi_{t}\mathrm{d}z$
$=2dm_{0}(\Phi_{t})+dm_{2}(\Phi_{t})+2m_{2}(\Phi_{t})+2\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z$
$~{}~{}+4\int\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),z\rangle\Phi_{t}\mathrm{d}z\,.$
The first term $\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z$ is upper-bounded by $\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}dm_{2}(\Phi_{t})$ and for the second term we have
$\int\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),z\rangle\Phi_{t}\mathrm{d}z$ $=\int\langle\nabla\log\left(\frac{\pi}{p_{t}^{V}}\right)(z)-\nabla\log\left(\frac{\pi}{p_{t}^{V}}\right)(0),z\rangle\Phi_{t}\mathrm{d}z$
$~{}~{}~{}+\int\langle\nabla\log\left(\frac{\pi}{p_{t}^{V}}\right)(0),z\rangle\Phi_{t}\mathrm{d}z$
$\leq\int\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}\|z\|^{2}\Phi_{t}\mathrm{d}z+\|\log(p_{t}^{V})(0)\|\int\|z\|\Phi_{t}\mathrm{d}z$
$=\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}m_{2}(\Phi_{t})+\|\nabla\log(p_{t}^{V})(0)\|m_{1}(\Phi_{t})\,.$
Hence we recover
$\partial m_{2}(\Phi_{t})\leq(d+2)\left(1+\frac{2(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}\right)m_{2}(\Phi_{t})+4\|\nabla\log(p_{t}^{V})(0)\|m_{1}(\Phi_{t})+2dm_{0}(\Phi_{t})\,.$
We now use the Gronwall lemma to obtain
$m_{2}(\Phi_{t})\leq e^{t(d+2)}(L(1-e^{-2t})+e^{-2t})^{d+2}\left(m_{2}(\Phi_{0})+2\int_{0}^{t}\frac{e^{-s(d+2)}(2m_{1}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|+dm_{0}(\Phi_{s}))}{(L(1-e^{-2s})+e^{-2s})^{(d+2)}}\mathrm{d}s\right)$
Recalling the upper-bound $m_{1}(\Phi_{t})\leq e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}C$ where $C$ is defined in Lemma 24, we upper-bound the integral term as
$\int_{0}^{t}\frac{e^{-s(d+2)}m_{1}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|}{(L(1-e^{-2s})+e^{-2s})^{(d+2)}}\mathrm{d}s$ $\leq C\int_{0}^{t}\frac{e^{-s}}{L(1-e^{-2s})+e^{-2s})}\|\nabla\log(p_{s}^{V})(0)\|\mathrm{d}s$
$\leq C\sqrt{\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s}\sqrt{\int_{0}^{t}\|\nabla\log(p_{t}^{V})(0)\|^{2}\mathrm{d}s}$
$=C\sqrt{\frac{1-e^{-2t}}{2}}\sqrt{\int_{0}^{t}\|\nabla\log(p_{t}^{V})(0)\|^{2}\mathrm{d}s}\,.$
Similarly,
$\int_{0}^{t}\frac{e^{-s(d+2)}}{(L(1-e^{-2s})+e^{-2s})^{(d+2)}}m_{0}(\Phi_{s})\mathrm{d}s$ $\leq\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}$
$=\frac{1-e^{-2t}}{2}\,.$
Hence we recover that
$m_{2}(\Phi_{t})\leq e^{t(d+2)}(L(1-e^{-2t})+e^{-2t})^{d+2}\left(m_{2}(\Phi_{0})+2C\sqrt{-2\log\left(\mu(0)\right)+d\log(\frac{L}{2\pi})}+d\right)\,.$
Using the expression of $C$, we recover eventually that
$m_{2}(\Phi_{t})\leq\frac{2e^{t(d+2)}Z_{V}^{2}}{Z_{2V}}(L(1-e^{-2t})+e^{-2t})^{d+2}\left[m_{2}(\mu)+d(L+1)-4\log\left(\mu(0)\right)+2d\log(\frac{L}{2\pi})\right]\,.$
A.7 Integrated average squared error of the estimator
Proposition 25
Under Assumptions 1-2 and the additional assumptions that $\|\nabla\log(p_{t}^{V})(0)\|,\|\nabla\log(p_{t}^{2V})(0)\|\leq\xi$ for some $\xi\geq 0$ and that $\mu$ has finite second moment $m_{2}(\mu)$, it holds that
$\begin{split}\int\mathbb{E}[\|\hat{s}_{n,t}(z)-\nabla\log(p_{t})(z)\|^{2}]\mathrm{d}p_{t}^{V}(z)\lesssim~{}&\frac{e^{2t(d+1)}(d+\log(n))(L+1)^{d}}{n(1-e^{-2t})^{2}}(K_{1}+K_{2})\,,\end{split}$
where we defined
$\begin{cases}&K_{1}=(L+\beta+1)^{2}(L+1)^{2}\bigg{[}m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\bigg{]}\,,\\ &K_{2}=d\beta+\xi^{2}+1\,.\end{cases}$ (10)
Proof.
Assumption 1 yields
$\Delta\log(p_{t}^{2V})\leq\beta d\,,$
and Assumptions 1-2 together with Proposition 9 yield
$\|\nabla^{2}\log(p_{t}^{iV})\|\leq(iL+\beta+1)\,,$
for $i\in\{1,2\}$. These estimates together with Proposition 7 yield for the integrated error:
$\begin{split}\int\delta(z)^{2}\mathrm{d}p_{t}^{V}(z)\lesssim&~{}e^{td}\left[d(\beta+\frac{1}{1-e^{-2t}})+\xi^{2}+1\right]\frac{(d+\log(n))Z_{2V}}{n(1-e^{-2t})(Z_{V})^{2}}\int\Phi_{t}(z)\mathrm{d}z\\ &+\frac{e^{td}Z_{2V}(d+\log(n))(L+\beta+1)^{2}}{n(Z_{V})^{2}(1-e^{-2t})}\int\|z\|^{2}\Phi_{t}(z)\mathrm{d}z\,.\end{split}$
We use Lemma 24 and we obtain
$\begin{split}\int\delta(z)^{2}\mathrm{d}p_{t}^{V}(z)\lesssim&~{}\frac{e^{2td}(d+\log(n))(L+1)^{d}}{n(1-e^{-2t})}\bigg{(}\left[d(\beta+\frac{1}{1-e^{-2t}})+\xi^{2}+1\right]\\ &+(L+\beta+1)^{2}(L+1)^{2}e^{2t}\bigg{[}m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\bigg{]}\bigg{)}\\ &\lesssim~{}\frac{e^{2t(d+1)}(d+\log(n))(L+1)^{d}}{n(1-e^{-2t})^{2}}\bigg{(}\left[d\beta+\xi^{2}+1\right]\\ &+(L+\beta+1)^{2}(L+1)^{2}\bigg{[}m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\bigg{]}\bigg{)}\,.\end{split}$
A.8 Proof of Theorem 11
Proof.
We run the recursion Eq. 3 with $T=\log((d+m_{2}(\mu))/\epsilon)$, $N=\frac{d^{2}(\log(L)+T)}{\epsilon}$, $t_{k}$ as indicated in Theorem 3, the approximate score $\hat{s}_{n_{k},t_{k}}(z)=\frac{-1}{(1-e^{-2t_{k}})}\frac{\sum_{i=1}^{n_{k}}y_{i}^{k}e^{-V((z-y_{i}^{k})e^{t_{k}}}}{\sum_{i=1}^{n_{k}}e^{-V((z-y_{i})e^{t_{k}}}}$ with $y_{i}^{k}\sim\mathcal{N}(0,(1-e^{-2t_{k}})I_{d})$ and
$n_{k}$ $=\frac{e^{2t_{k}(d+1)}d(L+1)^{d}}{\epsilon(1-e^{2t_{k}})^{2}}(K_{1}+K_{2})\coloneqq C\frac{e^{2t_{k}(d+1)}}{(1-e^{2t_{k}})^{2}}\,.$
By definition of $n_{k}$ and the other parameters, Theorem 3 yields
$\mathbb{E}[\text{KL}(\mu,\hat{p})]\lesssim\epsilon\,,$
where $\lesssim$ hides a universal constant and logarithmic factors in $\epsilon,L,\beta,d,m_{2}(\mu),\log(\mu(0))$. The total running time of the algorithm is given by
$\sum_{k=1}^{N}n_{k}$ $=C\sum_{k=1}^{N}\frac{e^{2t_{k}(d+1)}}{(1-e^{2t_{k}})^{2}}$
$\leq CN\frac{e^{2t_{N}(d+1)}}{(1-e^{2t_{1}})^{2}}$
$=\frac{CN}{(1-e^{2t_{1}})^{2}}(m_{2}(\mu)+d)^{2(d+1)}\epsilon^{-2(d+1)}$
$\lesssim\frac{CN}{t_{1}^{2}}(m_{2}(\mu)+d)^{2(d+1)}\epsilon^{-2(d+1)}\,.$
Using the recursion in Theorem 3, we have that $t_{1}=\frac{c}{L}$ where we recall $c=(\log(L)+T)/N$. This yields $t_{1}=\epsilon\log((m_{2}(\mu)+d)L/\epsilon)/(Ld^{2})$ and a total complexity of
$\sum_{i=1}^{N}n_{k}\lesssim d^{7}L^{d+2}\epsilon^{-2(d+3)}(m_{2}(\mu)+d)^{2(d+1)}[$ $(L+\beta)^{2}L^{2}(m_{2}(\mu)+dL+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right))$
$+(d\beta+\xi^{2})]\,.$
A.9 Sampling with unknown parameters
If there is no prior knowledge on $L$, we can use a constant step size at the cost of a quadratic factor on $L$.
Theorem 26 (Polynomial time sampling, unknown parameters)
Under the same assumptions as in Theorem 11, if we run algorithm Eq. 3 with constant step size, $T=\log(1/\epsilon)$, $N=1/\epsilon$ and with the stochastic score estimators $\hat{s}_{n_{k},t_{k}}$ with sample size $n_{k}=e^{2t_{k}(d+1)}/(1-e^{-2t_{k}})^{2}$, then, denoting $\hat{p}$ the stochastic distribution of the output $Y_{N}$, it holds that
$\begin{cases}&\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon(d+m_{2}(\mu)+d^{2}(\beta+L)^{2}+d(L+1)^{d}(K_{1}+K_{2}))\,,\\ &\sum_{i=1}^{N}n_{k}\leq\epsilon^{-2(d+5/2)}\log(1/\epsilon)^{2}\end{cases}$
,absent\,,,
where $\lesssim$ hides a universal constant and $poly(\log(1/\epsilon))$ factors.
Proof.
We simply use the second statement of Theorem 3 combined with the bound on the average integrated error in Proposition 25. ∎
Appendix B Proofs of Section 4
B.1 Proof of Proposition 14
Before proving the result, we need the following intermediate result.
Lemma 27
Let $V$ be an $L$-smooth, $\alpha$-strongly convex potential outside some ball centered $B_{R}(m)$. Then for any critical point $x^{*}$ of $V$, it holds
$\langle\nabla V(x),x-x^{*}\rangle\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
Proof.
For all $x\in\mathbb{R}^{d}$, the following identity holds
$\langle\nabla V(x),x-x^{*}\rangle=\int\langle x-x^{*},\nabla^{2}V(tx+(1-t)x^{*})(x-x^{*})\rangle\,.$ (11)
Using the equation above, we split the analysis in the case where the critical point $x^{*}$ belongs to $B_{R}(m)$ and in the case where it does not. In the first case, if $\|x-x^{*}\|\leq 2R$, we can coarsely lower bound as
$\langle\nabla V(x),x-x^{*}\rangle\geq-L\|x-x^{*}\|^{2}\,.$
In the case where $\|x-x^{*}\|>2R$, note that whenever $t\geq\frac{2R}{\|x-x^{*}\|}$, it holds that
$\|tx+(1-t)x^{*}-m\|$ $=\|x^{*}-m+t(x-x^{*})\|$
$\geq t\|x-x^{*}\|-\|x^{*}-m\|$
$>R\,,$
in other words, $tx+(1-t)x^{*}\not\in B_{R}(m)$. This allows to recover the lower bound
$\langle\nabla V(x),x-x^{*}\rangle$ $\geq-L\|x-x^{*}\|^{2}\frac{2R}{\|x-x^{*}\|}+\alpha\|x-x^{*}\|^{2}(1-\frac{2R}{\|x-x^{*}\|})$
$=\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
Hence we recover that
$\langle\nabla V(x),x-x^{*}\rangle$ $\geq-\mathds{1}_{x\in B_{R}(m)}L\|x-x^{*}\|^{2}+\mathds{1}_{x\not\in B_{R}(m)}(\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|)$
$\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
In the case where $x^{*}\not\in B_{R}(m)$, let us introduce the cone
$C=\{x\in\mathbb{R}^{d}~{}|~{}\exists t\in[0,1],tx+(1-t)x^{*}\in B_{R}(m)\}\,;$
note that we indeed have $B_{R}(m)\subset C$. By definition of $C$, if $x\not\in C$, it holds that
$\langle\nabla V(x),x-x^{*}\rangle\geq\alpha\|x-x^{*}\|^{2}\,.$
We now split the analysis in the case where $x\in C\cap B_{R}(m)$ and where $x\in C\cap\overline{B_{R}(m)}$. Denoting $x(t)=tx+(1-t)x^{*}$ we define $t_{1}=\inf\{t\in[0,1]~{}|~{}x(t)\in B_{R}(m)\}$ ; remark that $t_{1}=\frac{\|x(t_{1})-x^{*}\|}{\|x-x^{*}\|}$. By definition of $t_{1}$, it holds for any $t<t_{1}$ that $x(t)\not\in B_{R}(m)$. Thus, if $x\in C\cap B_{R}(m)$, we get the following lower-bound
$\langle\nabla V(x),x-x^{*}\rangle$ $\geq\alpha\|x-x^{*}\|^{2}\frac{\|x(t_{1})-x^{*}\|}{\|x-x^{*}\|}-L\|x-x^{*}\|^{2}\biggl{(}1-\frac{\|x(t_{1})-x^{*}\|}{\|x-x^{*}\|}\biggr{)}$
$\geq\alpha\|x-x^{*}\|^{2}\frac{\|x-x^{*}\|-\|x-x(1)\|}{\|x-x^{*}\|}-L\|x-x^{*}\|^{2}\biggl{(}1-\frac{\|x-x^{*}\|-\|x-x(1)\|}{\|x-x^{*}\|}\biggr{)}$
$\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
Now, if $x\in C$, note that $t<\frac{\|m-x^{*}\|-R}{\|x-x^{*}\|}$, then it holds that
$\|tx+(1-t)x^{*}-m\|$ $\geq\|x^{*}-m\|-t\|x-x^{*}\|$
$>R\,.$
Furthermore, introducing for $x\in\mathbb{R}^{d}$ the projection $p_{B}(x)$ onto the ball $B_{R}(m)$ we recall that $p_{B}(x)-x$ is (positively) colinear to $m-p_{B}(x)$. Thus it holds that
$\frac{\|m-x^{*}\|-R}{\|x-x^{*}\|}$ $=\frac{\|p_{B}(x^{*})-x^{*}\|+\|p_{B}(x^{*})-m\|-R}{\|x-x^{*}\|}$
$=\frac{\|p_{B}(x^{*})-x^{*}\|}{\|x-x^{*}\|}\,.$
Ne now prove that $\langle x-p_{B}(x^{*}),p_{B}(x^{*})-x^{*}\rangle\geq 0$. By the absurd, assume that $\langle x-p_{B}(x^{*}),p_{B}(x^{*})-x^{*}\rangle<0$, then
$\langle tx+(1-t)x^{*}-p_{B}(x^{*}),p_{B}(x^{*})-x^{*}\rangle<0\,.$
In particular, for all $t\in[0,1]$, it holds that
$\|tx+(1-t)x^{*}-m\|^{2}$ $=\|m-p_{B}(x^{*})+p_{B}(x^{*})-(tx+(1-t)x^{*})\|^{2}$
$=\|m-p_{B}(x^{*})\|^{2}+2\langle m-p_{B}(x^{*}),p_{B}(x^{*})-(tx+(1-t)x^{*})\rangle$
$~{}~{}~{}+\|p_{B}(x^{*})-(tx+(1-t)x^{*})\|^{2}$
$=\|m-p_{B}(x^{*})\|^{2}+\|p_{B}(x^{*})-(tx+(1-t)x^{*})\|^{2}$
$~{}~{}~{}+2\frac{\|m-p_{B}(x^{*})\|}{\|p_{B}(x^{*})-x^{*}\|}\langle p_{B}(x^{*})-x^{*},p_{B}(x^{*})-(tx+(1-t)x^{*})\rangle$
$>R^{2}\,,$
which contradicts $x\in C$. Hence $\langle tx+(1-t)x^{*}-p_{B}(x^{*}),p_{B}(x^{*})-x^{*}\rangle\geq 0$ which implies
$\|x-x^{*}\|^{2}$ $=\|x-p_{B}(x^{*})\|^{2}+2\langle x-p_{B}(x^{*}),p_{B}(x^{*})-x^{*}\rangle+\|p_{B}(x^{*})-x^{*}\|^{2}$
$\geq\|p_{B}(x^{*})-x^{*}\|^{2}\,.$
In particular, we are ensured that $\frac{\|p_{B}(x^{*})-x^{*}\|}{\|x-x^{*}\|}\leq 1$. We now split the analysis in the case where $x\in B_{R}(m)\cap C$ and where $x\in\overline{B_{R}(m)}\cap C$. In the first case, the previous computation implies the following lower-bound:
$\langle\nabla V(x),x-x^{*}\rangle$ $\geq\alpha\|x-x^{*}\|(\|m-x^{*}\|-R)-L(R+\|x-x^{*}\|-\|m-x^{*}\|)\|x-x^{*}\|$
$\geq\alpha\|x-x^{*}\|(\|x-x^{*}\|-\|x-m\|-R)$
$~{}~{}~{}-L(R+\|x-x^{*}\|-\|x-x^{*}\|+\|m-x\|)\|x-x^{*}\|$
$\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
Conversely, in the case where $x\in C\cap\overline{B_{R}(m)}$, define $t_{2}=\sup\{t\in[0,1]~{}|~{}x(t)\in B_{R}(m)\}$ ; again note that we have $t_{2}=\frac{\|x(t_{2})-x^{*}\|}{\|x-x^{*}\|}$. By definition, note that for any $t>t_{2}$, $x(t)\not\in B_{R}(m)$ hence we get the lower-bound
$\langle\nabla V(x),x-x^{*}\rangle$ $=\int_{0}^{t_{1}}\langle(x-x^{*}),\nabla^{2}V(tx+(1-t)x^{*})(x-x^{*})\rangle$
$~{}~{}~{}+\int_{t_{1}}^{t_{2}}\langle(x-x^{*}),\nabla^{2}V(tx+(1-t)x^{*})(x-x^{*})\rangle$
$~{}~{}~{}+\int_{t_{2}}^{1}\langle(x-x^{*}),\nabla^{2}V(tx+(1-t)x^{*})(x-x^{*})\rangle$
$\geq\alpha\|x-x^{*}\|^{2}\biggl{(}1-\frac{\|x(t_{2})-x^{*}\|}{\|x-x^{*}\|}+\frac{\|x(t_{1})-x^{*}\|}{\|x-x^{*}\|}\biggr{)}$
$~{}~{}~{}-L\|x-x^{*}\|(\|x(t_{2})-x^{*}\|-\|x(t_{1})-x^{*}\|)$
$\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
Hence we recover for all three cases $x\not\in C$, $x\in C\cap B_{R}(m)$ and $x\in C\cap\overline{B_{R}(m)}$ that
$\langle\nabla V(x),x-x^{*}\rangle\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
We can now prove our result.
Proof.
Using Lemma 27, we have the lower bound
$\langle\nabla V(x),x-x^{*}\rangle\geq\alpha\|x-x^{*}\|^{2}-2R(L+\alpha)\|x-x^{*}\|\,.$
Using the Laplacian identity for $f(x)=\|x-x^{*}\|^{2}$, we get that
$\int\langle\nabla V(x),x-x^{*}\rangle\mathrm{d}\mu(x)=d\,,$
which combined with the upper-bound above yields
$\int\alpha\|x-x^{*}\|^{2}\mathrm{d}\mu(x)\leq 2R(L+\alpha)\int\|x-x^{*}\|\mathrm{d}\mu(x)+d\,.$
Denoting $\theta^{2}=\mathbb{E}_{X\sim\mu}[\|X-x^{*}\|^{2}]$, it holds by Cauchy-Schwarz that
$\theta^{2}\leq\frac{2R(L+\alpha)}{\alpha}\theta+\frac{d}{\alpha}\,.$
This implies that $\theta^{2}$ is upper-bounded as
$\theta^{2}$ $\leq\frac{1}{2}\biggl{(}\frac{2R(L+\alpha)}{\alpha}+\sqrt{\frac{4R^{2}(L+\alpha)^{2}}{\alpha^{2}}+\frac{4d}{\alpha}}\biggr{)}^{2}$
$\leq\biggl{(}\frac{4R^{2}(L+\alpha)^{2}}{\alpha^{2}}+\frac{4R^{2}(L+\alpha)^{2}}{\alpha^{2}}+\frac{4d}{\alpha}\biggr{)}$
$=\frac{8R^{2}(L+\alpha)^{2}}{\alpha^{2}}+\frac{4d}{\alpha}\,.$
B.2 Proof of Corollary 16
Proof.
We start by using Proposition 13 which ensures that as long as $1-e^{-2t}\leq\frac{e^{-t}}{2L}$, it holds that
$\|\nabla^{2}\log(p_{t}^{V})\|\leq 2Le^{t}\,.$
In other terms, as long as
$e^{t}\leq\frac{1+\sqrt{16L^{2}+1}}{4L}\,,$
it holds that
$\|\nabla^{2}\log(p_{t}^{V})\|\leq\frac{1+\sqrt{1+16L^{2}}}{2}\,.$
Furthermore we recall that Proposition 9 yields the (uniform-in-time) lower-bound
$\nabla^{2}\log(p_{t}^{V})\succeq-(L+1)I_{d}\,.$
Finally, in the case where $1-e^{-2t}>\frac{e^{-t}}{2L}$, we have in particular that
$\frac{e^{-2t}}{(1-e^{-2t})^{2}}\leq 4L^{2}\,.$
Now recall the expression of the Hessian along the forward:
$\nabla^{2}\log(p_{t}^{V})(x)-\nabla^{2}\log(\pi)(x)=\frac{e^{-2t}}{1-e^{-2t}}\left(\frac{\text{Cov}_{q_{t,x}}(X)}{1-e^{-2t}}-I_{d}\right)\,,$
with $q_{t,x}\propto e^{-V(z)}e^{-\frac{e^{-2t}\|z-xe^{t}\|^{2}}{2}}$. Note that $\log(q_{t,x})$ has the same condition number as $V$. In particular, we can apply Proposition 14 combined with the previous bound and recover
$\nabla^{2}\log(p_{t}^{V})\preceq 4L^{2}\left[8R^{2}\left(\frac{L+2\alpha}{\alpha}\right)^{2}+\frac{4d}{\alpha}+1\right]I_{d}\,.$
Hence, it holds for all $t$ that
$\|\nabla^{2}\log(p_{t}^{V})\|\leq 4(L+1)^{2}\biggl{[}8R^{2}\biggl{(}\frac{L+2\alpha}{\alpha}\biggr{)}^{2}+\frac{4d}{\alpha}+1\biggr{]}\,.$
B.3 Remaining estimates
Proposition 28
Let $V$ be a $C^{2}$, $L$-smooth, $\alpha$-strongly convex potential outside some ball of radius $R$ such that $\nabla V(0)=0$. Then it holds that
$\begin{cases}&\|\nabla\log(p_{t}^{V})(0)\|^{2}\leq(L+1)^{2}m_{2}(\mu)\,,\\ &2d\log(L\mu(0)^{-2/d}/2\pi)\leq 4(R^{2}\alpha(\kappa+1)^{2}+\log(2\kappa))\,.\end{cases}$
We first prove the estimate on the gradient.
Proof.
Recall that for all $z\in\mathbb{R}^{d}$, the score of the current distribution $p_{t}$ reads
$\nabla\log(p_{t})(z)=\frac{-1}{(1-e^{-2t})}\frac{\int(z-xe^{-t})e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}\mathrm{d}x}{\int e^{-\frac{\|z-xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}\mathrm{d}x}\,.$
In particular, at $z=0$, the score can be re-written as an expectation
$\nabla\log(p_{t})(0)=\frac{e^{-t}}{1-e^{-2t}}\mathbb{E}_{q_{t}}[X]\,,$
with $q_{t}\propto e^{-\frac{\|xe^{-t}\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}$. Note that $q_{t}$ is also $\alpha$-strongly log-concave outside some ball of radius $R$, has condition number at most $\kappa$ and is such that $\nabla\log(q_{t})(0)=0$. Thus, Proposition 14 applies and we recover by Jensen inequality
$\|\nabla\log(p_{t})(0)\|^{2}\leq\frac{e^{-2t}}{(1-e^{-2t})^{2}}(8R^{2}(\kappa^{2}+1)+\frac{4d}{\alpha})\,.$
We now recall the alternative formula for the score
$\nabla\log(p_{t})(z)=\frac{-e^{t}}{(1-e^{-2t})}\frac{\int\nabla V((z-x)e^{t})e^{-V((z-x)e^{t})}e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}\mathrm{d}x}{\int e^{-V((z-x)e^{t})}e^{-\frac{\|x\|^{2}}{2(1-e^{-2t})}}\mathrm{d}x}\,.$
Making the change of variable $u=(z-x)e^{t}$ for both the numerator and the denominator, we recover
$\nabla\log(p_{t})(z)=\frac{-e^{t}}{(1-e^{-2t})}\frac{\int\nabla V(u)e^{-V(u)}e^{-\frac{\|z-e^{-t}u\|^{2}}{2(1-e^{-2t})}}\mathrm{d}u}{\int e^{-V(u)}e^{-\frac{\|z-e^{-t}u\|^{2}}{2(1-e^{-2t})}}\mathrm{d}u}\,.$
In particular, for $z=0$ we get
$\|\nabla\log(p_{t})(0)\|^{2}$ $\leq e^{2t}\mathbb{E}_{q_{t}}[\|\nabla V(X)\|]^{2}$
$\leq L^{2}e^{2t}(8R^{2}(\kappa^{2}+1)+\frac{4d}{\alpha})\,.$
Finally, remark that $t\mapsto\frac{e^{-t}}{1-e^{-2t}}$ decreases and $t\mapsto e^{t}$ increases. When we evaluate both functions at $t=\frac{1}{2}\log((L+1)/L)$ we recover in both cases $\sqrt{L}\sqrt{L+1}$ which implies that for all $t>0$, we have
$\|\nabla\log(p_{t}^{V})\|^{2}\leq(L+1)^{2}(8R^{2}(\kappa^{2}+1)+\frac{4d}{\alpha})\,.$
Before proving the second estimate, we need this intermediary result.
Lemma 29
Let $V$ be an $L$-smooth potential that is $\alpha$-strongly convex outside some ball of radius $R$ and such that $\nabla V(0)=0$. It holds that,
$V(x)-V(0)\geq\frac{\alpha}{2}\|x\|^{2}-R(L+\alpha)\|x\|\,.$
Proof.
We write the difference $V(x)-V(0)$ in integral form:
$V(x)-V(0)$ $=\int_{0}^{1}\langle\nabla V(tx),x\rangle\mathrm{d}t$
$=\int_{0}^{1}\frac{1}{t}\langle\nabla V(tx),tx\rangle\mathrm{d}t\,.$
Now we use Lemma 27 with $x^{*}=0$ to recover
$V(x)-V(0)$ $\geq\int_{0}^{1}\frac{1}{t}(\alpha t^{2}\|x\|^{2}-tR(L+\alpha)\|x\|)\mathrm{d}t$
$=\frac{\alpha}{2}\|x\|^{2}-R(L+\alpha)\|x\|\,.$
We can now prove the second estimate. ∎
Proof.
We recall that
$\frac{\mathrm{d}\mu(x)}{\mathrm{d}x}=\frac{e^{-V(x)}}{\int e^{-V(x)}\mathrm{d}x}\,,$
which can be re-written as
$\frac{\mathrm{d}\mu(x)}{\mathrm{d}x}=\frac{e^{-(V(x)-V(0))}}{\int e^{-(V(x)-V(0))}\mathrm{d}x}\,.$
By smoothness and since we assumed $\nabla V(0)=0$, the numerator can be lower-bounded by $e^{-L\|x\|^{2}}$. To upper bound the denominator, we use the previous result to recover
$\int e^{-(V(x)-V(0))}\mathrm{d}x$ $\leq\int e^{-\frac{\alpha}{2}\|x\|^{2}+R(L+\alpha)\|x|}\mathrm{d}x$
$=\int e^{-\frac{\alpha}{4}\|x\|^{2}-\frac{\alpha}{4}\|x\|^{2}+R(L+\alpha)\|x\|}\,.$
Now observe that we have for all $x$
$-\frac{\alpha}{4}\|x\|^{2}+R(L+\alpha)\|x\|\leq\frac{R^{2}(L+\alpha)^{2}}{\alpha}\,,$
which implies
$\int e^{-(V(x)-V(0))}\mathrm{d}x$ $\leq\int e^{\frac{R^{2}(L+\alpha)^{2}}{\alpha}}e^{-\frac{\alpha}{4}\|x\|^{2}}\mathrm{d}x$
$=e^{\frac{R^{2}(L+\alpha)^{2}}{\alpha}}\biggl{(}\frac{4\pi}{\alpha}\biggr{)}^{d/2}\,.$
Hence we finally recover
$\frac{\mathrm{d}\mu(x)}{\mathrm{d}x}\geq e^{-L\|x\|^{2}}e^{-\frac{R^{2}(L+\alpha)^{2}}{\alpha}}\biggl{(}\frac{4\pi}{\alpha}\biggr{)}^{-d/2}\,.$
B.4 Proof of Corollary 18
Proof.
The estimates in Proposition 28 and in Corollary 16 yield
$K_{1}\lesssim L^{4}(R^{4}\kappa^{4}+\frac{d^{2}}{\alpha^{2}})L^{2}(R^{2}\kappa^{2}+d(\frac{1}{\alpha}+L))\asymp L^{6}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{3}\,,$
and
$K_{2}\lesssim L^{2}(R^{2}\kappa^{2}+\frac{d}{\alpha})+dL^{2}(R^{2}\kappa^{2}+\frac{d}{\alpha})\asymp dL^{2}(R^{2}\kappa^{2}+\frac{d}{\alpha})\,.$
Hence we obtain that $K_{1}+K_{2}\asymp L^{6}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{3}$ and yields as a total complexity
$\sum_{k=1}^{N}n_{k}\lesssim d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)}\,.$
Appendix C Discussion on concurrent work
C.1 Tail matching IS
In Chak, [8], the author provides polynomial time guarantees via importance sampling for the problem of sampling from a distribution that is $L$-log smooth, $\alpha$-strongly log-concave outside a ball radius of $R$ centered at 00. The proposal $\nu$ is designed as follows: given some threshold $M$, the log density of $\nu$ reads $\log(\nu)(x)=-V(x)$ if $V(x)\geq M$ and $\log(\nu)(x)=M$ if $V(x)>M$. In Section 1, the author claims that if $M$ is chosen as $M=V(0)+\frac{LR^{2}}{2}$ (Equation (1.4)), $\nu$ can be safely sampled from using any standard MCMC algorithm such as Langevin. Yet, in the case for instance where $\mu=0.5\mathcal{N}(0,1)+0.5\mathcal{N}(m,1)$, the value of $R$ depends on $m$, and therefore the localization of the modes. Indeed, the second derivative of $V$ in this case is given by
$V^{\prime\prime}(x)=1-\frac{x^{2}e^{-x^{2}/2}+(x-m)^{2}e^{-(x-m)^{2}/2}}{e^{-x^{2}/2}+e^{-(x-m)^{2}/2}}+\frac{(xe^{-x^{2}/2}+(x-m)e^{-(x-m)^{2}/2})^{2}}{(e^{-x^{2}/2}+e^{-(x-m)^{2}/2})^{2}}\,.$
In particular, for $x=m/2$, it holds that
$V^{\prime\prime}(m/2)=1-(m/2)^{2}\,.$
Hence, whenever $|m|>2$, $V^{\prime\prime}(m/2)<0$ which implies $R\geq|m|/2$. We illustrate on Figure 1 the effect of a poorly chosen threshold $M$ in this particular example.
Figure 1: Target $\mu$ vs proposal $\nu$: the proposal remains multimodal hence cannot be easily sampled from with standard MCMC algorithms.
C.2 Self-normalized estimator for Follmer flow
In Ding et al., [18], the authors assume that there exists a known Gaussian distribution $\pi^{\eta,\Sigma}\sim\mathcal{N}(\eta,\Sigma)$ such that $\mathrm{d}\mu/\mathrm{d}\pi^{\eta,\Sigma}$ is (i) strictly bounded from above and (ii) is uniformly Lipschitz. In the simple case where $\mu$ is a Gaussian those two conditions imply that necessarily $\pi^{\eta,\Sigma}=\mu$. In the case where $\mu=0.5\mathcal{N}(m,1)+0.5\mathcal{N}(-m,1)$, condition (i) implies $\Sigma\geq 1$ and condition (ii) implies $\Sigma\leq 1$ which yields $\Sigma=1$. The ratio thus becomes
$\frac{\mathrm{d}\mu}{\mathrm{d}\pi^{\beta,\Sigma}}=e^{(\eta^{2}-m^{2})/2}\frac{e^{xm/2}+e^{-xm/2}}{e^{-\eta x/2}}\,.$
In particular, there exists no value of $\eta$ such that the ratio is uniformly Lipschitz.
Appendix D 2D simulation
We seek to sample from the 2D Gaussian mixture $\mu=0.5\mathcal{N}((0,0),2I_{2})+0.5\mathcal{N}((10,10),I_{2})$. We simulate $1000$ samples using ULA with step size $h=0.01$ and number of steps $N=1000000$ and using our algorithm with the same (uniform) step size $h$, $N=500$ steps and $1000$ samples to approximate the scores $\hat{s}_{n,t}$. Note that for this choice for parameters, the total running time of ULA was larger than the running time of our algorithm (on the same computer). We display on Figure 2 true samples vs samples obtained with ULA vs samples obtained with our algorithm. While ULA remained stuck in the mode $(0,0)$, our method accurately recovered the second mode $(10,10)$ and with the right proportion: denoting $\hat{w}$ the number of empirical samples $(x_{i})$ such that $\|(10,10)-x_{i}\|<5$, we found $\hat{w}\approx 0.48$.
Figure 2: Ground truth vs ULA vs our algorithm: ULA did not recover the top right mode whereas our algorithm did.